### v1

# mO Revanth[#1]

Abstract
The growing collaboration between humans and AI models in generative tasks has introduced new challenges in distinguishing between human-written, AI-generated, and human-AI collaborative texts. In this work, we collect a multilingual, multi-domain, multi-generator dataset FAIDSet. We further introduce a fine-grained detection framework FAID to classify text into these three categories, meanwhile identifying the underlying AI model family. Unlike existing binary classifiers, FAID is built to capture both authorship and model-specific characteristics. Our method combines multi-level contrastive learning with multi-task auxiliary classification to learn subtle stylistic cues. By modeling AI families as distinct stylistic entities, FAID offers improved interpretability. We incorporate an adaptation to address distributional shifts without retraining for unseen data. Experimental results demonstrate that FAID outperforms several baseline approaches, particularly enhancing the generalization accuracy on unseen domains and new AI models. It provide a potential solution for improving transparency and accountability in AI-assisted writing.
1 Introduction
AI has progressively evolved from an assistant tool to a creator or initiator, from helping polish papers and generate code to initiate proposals and draft essays, while humans increasingly serve as optimizers and reviewers. In such deeply collaborative human-AI settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al. (2025). This work aims to identify human involvement by distinguishing the origin of a given text: fully AI-generated, fully human-written, or collaboratively produced, in multilingual context.
Numerous studies have explored multilingual AI-generated text detection, but most focus either on binary detection: human vs. AI Wang et al. (2024d, b), or on fine-grained detection limited to English Abassy et al. (2024); Koike et al. (2024); Wang et al. (2023); Zhang et al. (2024a). Moreover, both struggle with generalization to unseen domains, languages, and new AI generators Wang et al. (2024c); Li et al. (2024).
Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain and multi-generator dataset for fine-grained detection, i.e., identifying a text into three categories: AI-generated, human-written, and human-AI collaborative; and (ii) introducing a detection framework FAID to ameliorate poor generalization performance. [^1] Our code and data are accessible at github.com/ngocminhta/FAID.
Our dataset focuses on academic field including domains of research paper abstracts, student theses, and scholarly reports, and involve recent LLM families, e.g., GPT, Gemini, DeepSeek and Llama OpenAI (2024); Team et al. (2024); DeepSeek-AI (2024); Dubey et al. (2024).
FAID treats each AI model as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features. This was achieved by optimizing a language encoder with multi-level author relationship (e.g., stylistic similarity between texts from the same LLM family is greater than the similarity between a human and an AI) to capture author-specific distinguishable signals of an input text based on contrastive learning, along with the task of fine-tuning a classifier to recognize its origin. This multitask learning process forces the encoder to reorganize the hidden space, representations of texts from the same authors distributed closer than from different authors.
Practically, given that generations produced by models developed by the same company tend to have similar writing styles due to similar model architecture, training data and training strategy (more details in Appendix C), we consider each AI family as an author. This can also help the detector acquire prior knowledge of future LLMs from the same company. Experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations. Our contributions can be summarized as follows:
  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained AI-generated text detection with 84k examples.
  • We propose a detection framework FAID to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in hidden representation.
  • Experimental results show that FAID outperforms baseline detectors, particularly over unseen domains and generators. Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to others with ground truth authorship in the database, implicitly enhancing interpretability with cues explaining why a text is assigned to a specific class.
2 Backgrounds
Advanced LLMs have changed how we create and refine text. Instead of originating every word, humans now more often act as post-editors or reviewers, stepping in after AI drafts the initial content. This collaborative workflow spans fields from academic publishing and journalism to education and social media, making hybrid human-AI texts the new norm Cheng et al. (2024); Lee et al. (2022).
Such a shift calls question into binary authorship detection: human vs. AI. Texts can be fully human-written, fully AI-generated, or collaborative text (e.g., human-written, AI-polished; human-initiated, AI-continued; AI-generated, human-edited; and deeply mixed text) Artemova et al. (2025). This demands the fine-granularity authorship detection, and even tracing back to a specific model family, to measure human contributions Hutson (2025). Then, clear attribution of AI involvement helps uphold honesty in research and preserve scientific integrity. Responding to these challenges, we collect a new dataset, and propose a detection framework which generalizes well to new domains, languages, and models, showing high accuracy and reliability.
2.1 Fine-grained AI-generated Text Datasets
Many prior studies have explored fine-grained AI-generated text datasets across various forms of human-AI collaboration as shown in Table 3, for example, MixSet (Zhang et al., 2024b), and Beemo (Artemova et al., 2025). More datasets and detailed information regarding the label space, and the coverage of domains, languages and LLMs for each are summarized in Table 3.
Despite these contributions, they have been limited to English. There is a noticeable gap in the availability of large-scale, fine-grained multilingual AI-generated text detection datasets. To bridge this gap, we collect a multilingual fine-grained AI-generated text detection dataset. It encompasses 84k texts generated by the latest LLMs and includes diverse forms of human-AI collaborative generations. This dataset can facilitate the development of more robust and generalizable detection models that are capable of handling complex multilingual collaborative scenarios.
2.2 Generalization of AI-generated Text Detection
Recent studies such as M4GT Wang et al. (2024c) and LLM-DetectAIve Abassy et al. (2024) have highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators. Many detection methods show a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors to deal with real-world scenarios and evolving LLMs outputs.
Various approaches have been proposed to improve generalization. OUTFOX Koike et al. (2024) leverages adversarial in-context learning, dynamically generating challenging examples to enhance robustness, but still faces limitations in domain transferability and computational efficiency. LLM-DetectAIve Abassy et al. (2024) adopts fine-grained multi-way classification and incorporates domain-adversarial training to reduce overfitting, yet its generalization to truly unseen domains and generators remains limited. Also, its current version lacks multilingual support. SeqXGPT Wang et al. (2023) focused on sentence-level detection by combining log-probability features with convolutional and self-attention mechanisms, which helps capture subtle mixed-content signals and improves generalization across input styles. However, its reliance on specific model features and limited semantic representation constrains adaptability to new generators and domains.
A recent representative work, the DeTeCtive framework Guo et al. (2024), introduces multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, though it primarily focuses on binary classification and does not fully address human-AI collaborative texts.
2.3 Contrastive Learning for AI-Generated Text Detection
Contrastive learning has been extensively used to improve the sentence representation by pulling semantically similar sentences closer and pushing dissimilar ones apart, reorganizing the hidden space of sentence representation by semantics. For example, a representative work SimCSE regarded an input sentence under dropout noise as its semantically similar counterpart (i.e., positive pairs), and trained the encoder to close their distance. Additionally, they natural language inference (NLI) pairs, treating entailment pairs as positives and contradiction pairs as hard negatives (Gao et al., 2021). For negative pairs with contradictory meanings, the embeddings were trained to be distant in the latent space. Similarly, DeCLUTR Giorgi et al. (2021) constructed positive pairs by extracting different spans from the same texts, and negative pairs are sampled from different texts.
We adopt this philosophy to reorganize the latent space based on writing styles, pulling human-written texts to cluster together while remaining distant from AI-generated ones. Analogous to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement in collaborative texts, ranging from 0 (fully AI-generated) to 1 (fully human-written).
DeTeCtive also took advantage of contrastive learning and was used in binary AI-generated text detection Guo et al. (2024). Based on a multi-task framework, DeTeCtive was trained to learn writing-style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs. AI) to capture distinguishable signals. For inference, the pipeline encoded the input text as a hidden vector and applied dense retrieval to match the cluster, depending on the stylistic similarity with previously indexed training feature database. Additionally, instead of retraining the model when faced with new data, they encoded new data with the trained encoder to obtain embeddings, and then added them to the feature database to augment. This largely improved the generalizability in unseen domains and new generators.
However, this approach only considers two types of text: human-written and AI-generated, without further consideration of human-AI collaborative texts. Our work fills this gap with the goal of improving generalization performance for fine-grained AI-generated text detection.
3 FAIDSet
To fill the gap of fewer publicly-available multilingual fine-grained AI-generated text datasets, we collected a multilingual, multi-domain and multi-generator AI-generated text dataset FAIDSet. It contains texts generated by LLMs, written by humans and collaborated by both humans and LLMs, resulting in a total of $\sim$84k examples. See Appendix B for more details about FAIDSet.
FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages including Vietnamese and English. We collected student theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^2] https://vjol.info.vn/ (VJOL).
Models and Label Space. We used the following multilingual LLMs families to produce both AI and human-AI collaborative texts: GPT-4/4o, Llama-3.x, Gemini 2.x, and Deepseek V3/R1. Regarding the human-AI collaborative text, we include AI-polished, AI-continued, and AI-paraphrased where the models are requested to polish or paraphrase inputs while ensuring the accuracy of any figures and statistics.
Diverse Prompt Strategies. We generate data with diverse tones and contexts while ensuring the accuracy of the content. Depending on the data source and context, we craft prompts to create varied outputs suitable for different real-world application scenarios. We generate responses with different tones using prompts such as "You are an IT student…" and "…who are very familiar with abstract writing…", which are described more in Appendix B.2.
Quality Control. Given that the task of AI-generated text detection naturally has gold labels when collecting data, we perform quality control by randomly sampling 10-20 instances for each domain, source, and generative model in data collection, and manually assess whether or not the generated texts are fluent and logically reasonable, to avoid incoherence, disorganization, or illogical texts. Such cases introduce remarkably distinguishable signals from human-written text, affecting the data quality. We adjust either the prompts or the generation parameters to address these problems.
4 Methodology
In this section, we first clarify task definition and then elaborate our detection method.
4.1 Task Definition
In this work, we formulate the task as a three-class classification problem: human-written, AI-generated, and human-AI collaborative texts. The human-AI collaborative category involves a range of interaction between humans and AI systems, such as human-written, AI-polished; human-initiated, AI-continued; human-written, AI-paraphrased; AI-generated, AI-humanized, etc.
Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, we do not exhaust all patterns. Instead, we consolidate all forms of human-AI collaboration into a single label to maintain practical simplicity and model generalizability. This approach reflects real-world usage, where using AI tools to enhance clarity or expression is increasingly common and often ethically acceptable.
We also investigate the dataset and find that the AI models within the same family tend to have similar writing styles and text distributions, due to their shared training data and architecture (see Appendix C). Therefore, we consider each model’s family to be an author with unique writing styles.
4.2 Framework Overview
Figure 1: Training architecture. Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix D) based on the human, human-AI and AI texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther. We train an encoder that can represent text with distinguishable signals to discern authorship of text.
We aim to enable the encoder to capture multi-level similarities between authors (AI, human and human-AI) as Eq (6), forcing model to learn features that can differentiate distinct authors in a high-dimensional hidden vector space. We denote $\mathrm{S}_{c}$ as cosine similarity, $\phi(\cdot)$ is the encoder function and $P_{i}$, $P_{j}$, ($51\leq i\leq j<$ 5) are distributions of different text subsets. We have:
$\begin{split}\mathbb{E}_{x\in P_{i},y\in P_{j}}&\left[\mathrm{S}_{c}\left(\phi(x),\phi(y)\right)\right]\\ &\geq\mathbb{E}_{x\in P_{i},y\in P_{j+1}}\left[\mathrm{S}_{c}\left(\phi(x),\phi(y)\right)\right]\end{split}$ (1)
where $P_{1}$ corresponds to the distribution generated by a particular LLM family, $P_{2}$ to the distribution generated by any LLMs, $P_{3}$ to the distribution of a collaborative text generated by humans and the LLM family of $P_{1}$, $P_{4}$ to the distribution of collaborative text generated by humans and any LLM families, and $P_{5}$ to the distribution of human-written text.
To clarify the rationale for configuring FAID to expect that the similarity of a text $x$ (from lower-level distributions $P_{1}$ or $P_{2}$) with samples from $P_{3}$ is generally greater than or equal to its similarity with samples from $P_{4}$, consider the following.
  • If $x\in P_{1}$ ($x$ is generated by a particular LLM): Let $y_{\textit{LHS}}$ be drawn from $P_{3}$ and $y_{\textit{RHS}}$ be drawn from $P_{4}$. Naturally, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ should be greater than $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. This is because $P_{3}$ contains texts that share a direct LLM family origin with $x$.
  • If $x\in P_{2}$ ($x$ is generated by any LLM): In this scenario, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$. Since $x$ can originate from any LLM, it does not inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
This configuration aims to ensure that closeness in distribution corresponds to heightened similarity after encoding, encouraging the model to discern fine-grained multi-level relations.
4.3 Multi-level Contrastive Learning
Given a dataset with $N$ samples, the $i^{th}$ sample is denoted as $T_{i}$ and assigned with three-level labels indicating its source.
  • $x$: if $T_{i}$ is fully AI-generated, $x_{i}=0$, otherwise $x_{i}=1$;
  • $y$: if $T_{i}$ is fully human-written, $y_{i}=0$, otherwise $y_{i}=1$;
  • $z$: an indicator of a specific LLM family.
The encoder $\phi(\cdot)$ represents the sample $T_{i}$ in a $d$-dimensional vector space $\mathbb{R}^{d}$. Then we calculate the cosine similarity between two samples $T_{i}$ and $T_{j}$, denoted by: $\sigma(i,j)=\mathrm{S}_{c}\left(\phi(T_{i}),\phi(T_{j})\right)$.
For AI-generated text ($x=0$), the similarity between $T_{i}$ and another AI-generated text $T_{j}$ is greater than that with a human-written or collaborative text $T_{k}$ ($x=1$):
$\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,x_{i}=x_{j},x_{k}=1$ (2)
If $x=0$, then the text is fully AI-generated. For this case, we do not consider $y$ since AI-generated text is considered as non-AI-collaboration. We can imply that the similarity between two texts written by the same AI is higher than that of two different AI families. Hence:
$\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,z_{i}=z_{j},z_{i}\neq z_{k}$ (3)
The reverse condition (with $y=1$) is also true. We can conclude that:
$\sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i}=y_{j},y_{i}\neq y_{k}$ (4)
For cases where all samples are human-AI collaborative texts ($x,y=1$), two texts with involvements of the same AI family tend to be more similar than the ones that involve contributions from different families. That is:
$\sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i,j,k}=1,z_{i}=z_{j}\neq z_{k}$ (5)
Combining all, the text representation is learned with the following constraints:
$\begin{cases}\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,x_{i}=x_{j}\neq x_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=0,z_{i}=z_{j}\neq z_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=1,x_{i}=x_{j}\neq x_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i}=y_{j}\neq y_{k};\\ \begin{split}\sigma(i,j)>\sigma(i,k),\forall&x_{i}=1,y_{i,j,k}=1,\\ &z_{i}=z_{j},z_{i}\neq z_{k}\\ \end{split}\end{cases}$
To enforce the similarity constraints outlined in Eq (6), we build upon the SimCLR framework Chen et al. (2020) and introduce a strategy for defining both positive and negative sample pairs, which forms the basis of our contrastive learning loss. Departing from traditional contrastive losses that rely on a single positive sample, our approach considers a group of positive instances that satisfy specific criteria. The similarity between the anchor and positive samples is computed as the average similarity across this entire positive set. For negative samples, we follow the methodology used in SimCLR. The resulting contrastive loss, expressed in Eq (7), involves $q$ as the anchor sample, $K^{+}$ as the positive sample set, $K^{-}$ as the negative sample set, $\tau$ as the temperature parameter, and $N_{K^{+}}$ as the number of positive samples.
$\mathcal{L}_{q}=-\log\frac{\exp\left(\sum_{k\in K^{+}}\frac{S(q,k)}{\tau}/N_{K^{+}}\right)}{\exp\left(\sum_{k\in K^{+}}\frac{S(q,k)}{\tau}/N_{K^{+}}\right)+\sum_{k\in K^{-}}\exp\left(\frac{S(q,k)}{\tau}\right)}$
Different constraints result in varied sets of positive and negative samples. Based on these sets, contrastive losses are computed at multiple levels. As we declared in Eq (7), each inequality in Eq (6) is denoted as $\mathcal{L}_{q_{i},\varepsilon}$ where $\varepsilon=\overline{1,5}$ respectively. To form Eq (8), we need to add the coefficients $\alpha,\beta,\gamma,\delta$, and $\zeta$ to maintain the balance of multi-level relations.
$\begin{split}\mathcal{L}_{mcl}=&\sum\limits_{i=1}^{N}\left[x_{i}\left(\alpha\mathcal{L}_{q_{i},1}+\beta\mathcal{L}_{q_{i},2}\right)+\right.\\ &\left.\left(1-x_{i}\right)\left(\gamma\mathcal{L}_{q_{i},3}+\delta\mathcal{L}_{q_{i},4}+\zeta\mathcal{L}_{q_{i},5}\right)\right]\end{split}$ (8)
Therefore, due to the last inequality in Eq (6) only specifying a case ($y=1$), and the other cases considering both values for $y$, so we have $\zeta=2\gamma=2\delta$. Also, to maintain the equilibrium, we need to keep $\alpha+\beta=\gamma+\delta+\zeta$. We set $\gamma=\delta=1$, then $\zeta=2,\alpha=\beta=2$.
This approach encourages the model to capture subtle and detailed features from different sources. As a result, the model becomes more adept at recognizing variations in writing styles. This capability enhances its accuracy and strengthens its ability to generalize when detecting AI-generated text.
4.4 Multi-task Auxiliary Learning
Multi-task learning Caruana (1997) allows a model to learn several tasks concurrently by sharing relevant information across them. This joint learning process helps the model develop more general and distinctive features. As a result, it improves the model’s ability to generalize to new data. Building on the previously described contrastive learning framework, we extend the encoder by attaching an MLP classifier at its output layer. This classifier is responsible for binary classification, determining whether a given query text was written by a human or generated by a LLM. Denote that the probability of $i^{th}$ sample with label $x_{i}=0$ is $p_{i}$. To train this component, we apply a cross-entropy loss function, denoted as $\mathcal{L}_{ce}$, defined as follows:
$\mathcal{L}_{ce}=-\dfrac{1}{N}\sum\limits_{i=1}^{N}x_{i}\log p_{i}+\left(1-x_{i}\right)\log\left(1-p_{i}\right)$ (9)
Therefore, the overall loss is computed as:
$\mathcal{L}=\mathcal{L}_{ce}+\mathcal{L}_{mcl}$ (10)
Figure 2: Inference architecture. (a) embed the input text into embedding vector using the fine-tuned encoder; (b) use Fuzzy KNN algorithm to cluster, retrieving which cluster the input text belongs to (see more in Appendix E). (c) The stored vector database $\mathcal{VD}$ was created by saving all embeddings of texts in training and validation sets using the fine-tuned encoder. If the input text is unseen, we embed the unseen text and save into a temporary vector database $\mathcal{VD}^{\prime}$, enhancing the generalization of the detector.
4.5 Handling Unseen Data without Retraining
Unseen data, whether from an unseen domain or generated by an unfamiliar model, remains a significant challenge even for state-of-the-art AI-generated text detection methods, due to the increasing capabilities of large language models. We tried to use the model classifier on its own, but we ended up by using a vector database along with Fuzzy k-Nearest Neighbors, which is illustrated in Figure 2. The results were written in Appendix E. Specifically, when dealing with unseen data, we use our trained model to embed these samples and add them to our existing vector database. Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without the need retraining.
5 Experiments
In this section, we first describe datasets and baselines used, followed by two experiments evaluating FAID in classifying a text between human, AI and human-AI; and in identifying specific generators.
Scenario Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
In-domain, Known generator FAIDSet LLM-DetectAIve 94.34 94.45 93.79 94.10 0.1888 0.1107
SVM (N-grams) 84.41 84.53 82.93 83.62 0.5630 0.2916
T5-Sentinel 93.31 94.92 93.10 93.15 0.2104 0.1101
FAID 95.58 95.78 95.33 95.54 0.1719 0.0875
LLM-DetectAIve LLM-DetectAIve 95.71 95.78 95.72 95.71 0.1606 0.1314
SVM (N-grams) 83.43 76.09 71.84 73.60 0.2878 0.2064
T5-Sentinel 94.77 94.70 92.60 93.60 0.1663 0.1503
FAID 96.99 95.29 88.14 91.58 0.1561 0.0754
HART LLM-DetectAIve 94.39 94.25 94.33 94.29 0.3244 0.1789
SVM (N-grams) 58.93 59.75 61.07 60.30 1.2334 0.6849
T5-Sentinel 86.68 87.25 87.69 87.38 0.4339 0.2334
FAID 96.73 97.61 98.05 97.80 0.4631 0.1806
Unseen domain, Unseen generator Unseen domain LLM-DetectAIve 52.83 47.31 64.62 53.28 0.4733 0.4722
SVM (N-grams) 39.28 43.13 29.41 22.77 0.7689 0.6611
T5-Sentinel 55.56 49.54 66.67 55.34 0.4444 0.4444
FAID 62.78 70.73 71.77 69.46 0.4514 0.4486
Unseen generators LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
SVM (N-grams) 74.19 61.88 50.17 55.17 0.4324 0.3162
T5-Sentinel 85.95 85.77 84.59 85.16 0.3648 0.2419
FAID 93.31 92.40 94.44 93.25 0.1691 0.1167
Unseen domain and Unseen generators LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
SVM (N-grams) 34.50 39.38 26.67 21.70 0.9186 0.7429
T5-Sentinel 57.07 49.82 66.61 55.45 0.4314 0.4300
FAID 66.55 74.44 73.57 72.58 0.3939 0.3167
Table 1: Performance with three labels. For in-domain and known generator datasets, we train detectors on training set and evaluate using test set. For unseen data evaluation, we use the detector trained on FAIDSet, and evaluate on the unseen datasets. The best results are in bold and the second best are underlined.
5.1 Datasets
In addition to FAIDSet, for in-domain evaluation, we used another two datasets:
LLM-DetectAIve Abassy et al. (2024) encompasses various domains including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews. The original human-written and machine-generated texts were augmented using a variety of LLMs (e.g., Llama 3, Mixtral, Gemma) to create 236,000-example dataset with two new labels: machine-written then machine-humanized, and human-written then machine-polished.
HART Bao et al. (2025) is a dataset spans domains of student essays, arXiv abstracts, story writing, and news articles. It has four categories: human-written texts, LLMs-refined texts, AI-generated texts, and humanized AI-generated texts. They further generated 21,500 cases based on this dataset for unbalanced labels.
To evaluate generalization ability of FAID on unseen scenarios, we collected the following data:
Unseen domain: We create a dataset consisting of 150 IELTS Writing essays from Kaggle ibrahimmazlum (2023), where all original texts are human-written. These essays were then used to generate human-AI collaborative and AI-generated texts, using the same models with FAIDSet.
Unseen generators: We selected 150 human-written abstracts from the FAIDSet test set and generated data for remaining labels using the three new LLM families: Qwen 2, Mistral, and Gemma 2.
Unseen domain & generators: Based on the human-written IELTS essays above, instead of the original four LLMs, we used the same LLM families as Unseen generator test set to generate data for the AI and human-AI labels, following the same process as with the unseen domains.
Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$
FAIDSet LLM-DetectAIve 0.7596 0.7697 0.7690 0.7653
SVM (N-grams) 0.6764 0.6515 0.6126 0.6312
T5-Sentinel 0.7568 0.7985 0.7840 0.7837
FAID 0.7964 0.8328 0.8352 0.8327
LLM-DetectAIve LLM-DetectAIve 0.9049 0.9064 0.8352 0.8693
SVM (N-grams) 0.8567 0.6516 0.6127 0.6313
T5-Sentinel 0.8154 0.8137 0.8009 0.8105
FAID 0.9089 0.8817 0.8672 0.8737
HART LLM-DetectAIve 0.8900 0.8787 0.8674 0.8715
SVM (N-grams) 0.6347 0.5454 0.4385 0.4806
T5-Sentinel 0.7852 0.7713 0.7834 0.7759
FAID 0.8996 0.9157 0.8648 0.8667
Table 2: Accuracy of identifying generators: human, GPT, Gemini, Deepseek, and Llama. The best is in bold.
5.2 Baselines
LLM-DetectAIve:
Abassy et al. (2024) fine-tuned a RoBERTa Liu et al. (2019a) on fine-grained English AI-generated texts with learning rate of $2\times 10^{-5}$ for 10 epochs.
Support Vector Machine with N-grams:
Prova (2024) utilized SVM with N-gram features for binary AI text detection (Human vs. AI). We adapted this approach to a three-class detection. We first generated word-level trigram features using Count Vectorizer, and then used a Linear SVM to train with the gold labels to guide the classification.
T5-Sentinel:
Chen et al. (2023) utilized T5-Sentinel to classify text into two labels (Human vs. AI). This approach leverages the T5 model to predict the conditional probability of the next token, integrating the classification task into a sequence-to-sequence completion framework. This unique method directly leverages the strength of encoder-decoder model, which are trained to understand context and generate coherent text.
5.3 Human, AI or Human-AI?
Table 1 shows the performance of FAID and three baselines in four evaluation settings. FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domain, (iii) unseen generators, and (iv) unseen domain & generators settings, followed by LLM-DetectAIve in (i) and (iv), and T5-Sentinel in (ii) and (iii). Taking advantage of pretrained language models (i.e., RoBERTa, T5), they outperform SVM in most cases, particularly on HART dataset.
FAID’s approach enhances the generalization performance over unseen domains and generators, compared with other methods. From the absolute accuracy, we can find that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55% respectively. The result suggests that FAID is an effective method to address the multilingual fine-grained AI-generated text detection task, improving the performance by leveraging multi-level contrastive learning to capture generalizable stylistic differences tied to AI families, rather than overfitting to surface-level artifacts.
5.4 Identifying Different Generators
We further examine the detector to identify specific AI generators which were seen during training. The goal of FAID is not only to detect AI involvement but also to identify the specific AI model or family applied, treating them as distinct authors. As shown in Table 2, across the three datasets, FAID consistently achieves higher performance compared to other baselines in almost all metrics. LLM-DetectAIve achieves comparable results with our detector FAID on the test set of LLM-DetectAIve dataset, except that its precision is slightly lower. FAID’s high performance when dealing with text originating from diverse known generators within these datasets indicates that FAID learned unique writing patterns and features of different AI generators by leveraging multi-level contrastive learning.
6 Conclusion and Future Works
In this work, we collected a multilingual fine-grained AI-generated text detection dataset FAIDSet with $\sim$84,000 examples. We further introduced FAID, a fine-grained AI-generated text detection framework aiming to distinguish between human-written, AI-generated, and human-AI collaborative texts. It combines multi-level contrastive learning and multi-task auxiliary objectives, with the perspective of treating AI model families as stylistic authors. FAID captures subtle writing patterns that generalize effectively across domains and evolving AI systems, where Fuzzy KNN-based inference method and training-free incremental adaptation enhance the robustness, enabling strong performance on unseen data.
Experiments demonstrate that FAID consistently outperforms other strong baselines across multiple datasets and configurations. Its capability to recognize nuanced forms of collaborative texts and the ability to adapt to new generative models make it particularly suited for real-world applications, improving transparency in AI-assisted writing.
In the future, we plan to expand FAIDSet by incorporating more languages, generators and domains, especially low-resource languages and informal text genres such as social media content or student essays, to further enhance cross-lingual and domain robustness. We also continue to investigate to deal with AI-generated text or more complex human-AI collaborative texts (i.e., from two or more AI families, generated by multiple AI families then edited by human) to deal with human’s behaviors in using AI for writing.
7 Limitations
While FAID demonstrates strong performance and generalization across various domains and AI models, several limitations remain. First, although our dataset is multilingual and multi-domain, it is still limited in terms of low-resource languages and niche writing domains, which may affect performance in those contexts. Second, our framework is based on the fact that texts produced by AI models from the same family share similar stylistic features. However, this assumption may break down when a single text is influenced by multiple AI systems – such as when a human uses different models for drafting, rewriting, and polishing. In such cases, the resulting style may blend traits from multiple sources, making it more difficult to attribute authorship to a single AI family or clearly distinguish collaboration boundaries.
8 Ethics and Broader Impacts
Data Collection and Licenses
A primary ethical consideration is the data license. We reused existing dataset for our research: LLM-DetectAIve, HART and IELTS Writing, which have been publicly released under clear licensing agreements. We adhere to the intended usage of these dataset licenses.
Security Implications.
FAIDSet streamlines both the creation and rigorous testing of FAID. By spotting AI-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors. More broadly, it supports efforts to prevent the misuse of generative technologies in areas such as credential falsification. Detecting AI-crafted content across different languages can be tricky, due to language’s unique grammar and style. By enabling robust, multilingual and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly. At the same time, it fosters critical digital literacy, giving everyone a clear understanding of both the strengths and limits of generative AI.
Responsible Use of AI-generated Text Detection.
FAID is designed to enhance transparency in AI-assisted writing by enabling the fine-grained detection of AI involvement in text generation. While this has clear benefits for academic integrity, and content provenance, we acknowledge the potential for misuse. For instance, such tools could be used to unfairly penalize individuals in educational or professional settings based on incorrect or biased predictions. To mitigate this, we stress that FAID is not intended for high-stakes decision-making without human oversight.
Bias and Fairness.
AI-generated text detection systems may inadvertently encode or amplify biases present in training data. FAIDSet has been carefully constructed to include diverse domains and languages to reduce such biases. Nonetheless, we encourage ongoing auditing and benchmarking of fairness across populations and writing styles, and welcome community feedback for further improvements.
Transparency and Reproducibility.
In this work, to promote open research and community contributions, we published our code and data to the public.
Appendix
Appendix A Some current datasets on AI-generated texts
We do analysis in many current dataset on AI-generated text and summarized them in Table 3. Many dataset contains fine-grained labels and multi-generator, but all of analyzed datasets are monolingual – only in English. This problem makes us to collect and formulate a new dataset for multilingual, multi-domain and multilingual AI-generated text dataset to adapt the approach to real-world scenarios.
Dataset Languages Label Space Domains Generators Size
MixSet
Zhang et al. (2024c)
English
Human-written, AI-polished
Human-initiated, AI-continued
AI-written, human-edited
Deeply-mixed text
Email
News
Game reviews,
Paper abstracts,
Speech,
Blog
GPT-4
Llama 2
Dolly
3,600
DetectiAIve
Abassy et al. (2024)
English
Human-written
AI-generated
Human-written, AI-polished
AI-written, AI-humanized
arXiv abstracts,
Reddit posts,
Wikihow,
Wikipedia articles,
OUTFOX essays,
Peer reviews
GPT-4o
Mistral 7B
Llama 3.1 8B
Llama 3.1 70B
Gemini
Cohere
487,996
Beemo
Artemova et al. (2025)
English
AI-generated, AI-humanized
Human-written
AI-generated
AI-written, human-edited
Generation,
Rewrite,
Open QA,
Summarize,
Closed QA
Llama 2
Llama 3.1
GPT-4o
Zephyr
Mixtral
Tulu
Gemma
Mistral
19,256
M4GT
Wang et al. (2024c)
English
Human-written, AI-continued
Human-written
AI-generated
Peer review,
OUTFOX
Llama 2
GPT-4
GPT-3.5
33,912
Real or Fake
Dugan et al. (2022)
English
Human-written
Human-initiated, AI-continued
Recipes,
Presidential Speeches,
Short Stories,
New York Times
GPT-2,
GPT-2 XL
CTRL
9,148
RoFT-chatgpt
Kushnareva et al. (2024)
English Human-initiated, AI-continued
Short Stories,
Recipes,
New York Times,
Presidential Speeches
GPT-3.5-turbo 6,940
Co-author
Zeng et al. (2024)
English Deeply-mixed text
Creative writing,
New York Times
GPT-3 1,447
TriBERT
Zeng et al. (2023)
English
Human-initiated, AI-continued
Deeply-mixed text
Human-written
Essays ChatGPT 34,272
LAMP
Chakrabarty et al. (2025)
English AI-generated, human-edited Creative writing
GPT-4o
Claude 3.5 Sonnet
Llama 3.1 70B
1,282
APT-Eval
Saha and Feizi (2025)
English Human-written, AI-polished Based on MixSet
GPT-4o
Llama 3.1 70B
Llama 3.1 8B
Llama 2 7B
11,700
HART
Bao et al. (2025)
English
Human-written
Human-written, AI-polished
AI-generated, AI-humanized
AI-generated text
AI-generated, human-edted
GPT-3.5-turbo
GPT-4o
Claude 3.5 Sonnet
Gemini 1.5 Pro
Llama 3.3 70B
Qwen 2.5 72B
16,000
LLMDetect
Cheng et al. (2024)
English
Human-written
Human-written, AI-polished
Human-written, AI-extended
AI-generated
DeepSeek v2
Llama 3 70B
Claude 3.5 Sonnet
GPT-4o
64,304
ICNALE corpus
Macko et al. (2025)
English Human-written Essays
Qwen 2.5
Llama 3.1 8B/70B
Llama 3.2 1B/3B
Mistral Small
67,000
CyberHumanAI
Najjar et al. (2025)
English Human-written Wikipedia 500
Table 3: English Fine-grained AI-generated Text Detection Datasets Overview.
Appendix B FAIDSet Statistics and Analysis
B.1 Statistics
Our FAIDSet includes 83,350 examples, which are separated into three subsets: train, validation, and test with the ratio shown in Table 4. The data set also comprises various sources of human-written texts, which are described in Table 5.
Subset Human AI Human-AI collab.
Train 14,176 12,076 32,091
Validation 3,038 2,588 6,876
Test 6,879 2,599 3,038
Table 4: Number of examples per label in subsets in FAIDSet dataset.
Source No. Human texts
arXiv abstracts 2,000
VJOL abstracts 2,195
HUST theses (English) 4,898
HUST theses (Vietnamese) 11,164
Table 5: Statistics of human-written text’s origins in FAIDSet.
B.2 Diverse Prompt Strategies
To avoid biasing our generated corpora toward any single style or topic, we employ a broad set of prompt templates when synthesizing AI-generated and human-AI collaborative texts. By varying prompt structure, content domain, and complexity, we ensure that the resulting outputs cover a wide distribution of writing patterns, vocabulary, and rhetorical devices. This diversity helps our detector generalize more effectively to real-world inputs.
Concretely, we sample two out of five prompts for AI-generated texts and several categories of human-AI collaborative texts in Table 6:
  • AI-polished: Texts that were originally written by a human and then lightly refined by an AI system to improve grammar, clarity, or fluency without altering the core content or intent.
  • AI-continued: Texts where a human wrote an initial portion (e.g., a sentence or paragraph), and an AI generated a continuation that attempts to follow the original style, tone, and intent.
  • AI-paraphrased: Texts that were originally written by a human and then reworded by an AI system to express the same meaning using different phrasing, possibly altering sentence structure or word choice while preserving the original message.
By mixing prompts across these categories, we generate a balanced corpus that mitigates over-fitting to any one prompt pattern and better reflects the diversity of real user queries.
Category Student Thesis Paper Abstract
AI-generated You are an IT student who are writing the graduation thesis. In a formal academic tone, write this section of a thesis on       , clearly outlining the structure, then write the passage concisely. The original text:       . In clear, structured prose, draft the section for a thesis titled       , cite some related works you mentioned in the passage and highlighting the contribution. The original text:       . You are a computer scientist who are very familiar with abstract writing for your works based on the title. Craft a concise word_count-word abstract for a paper titled       , summarizing the problem statement, methodology, key findings, and contributions. The original text:       . Compose a word_count-word abstract for the paper       , ensuring it includes motivation, approach, results, and implications for future research. The original text:       .
AI-polished You are an IT student who are refining your work to make it more complete. Refine the following section excerpt for grammar, clarity, and academic style while preserving its original meaning and terminology. The original text:       . You are an IT student who are refining your work to make it more complete. Enhance the academic tone, coherence, and logical flow of this thesis section without altering technical content. The original text:       . You are a computer scientist who are very familiar with abstract writing and refining the written abstract. Improve the coherence, precision, and formal tone of this draft abstract without introducing new content. The original text:       . Improve the clarity and conciseness of this abstract paragraph while maintaining all original findings and terminology. The original text:       .
AI-continued You are an IT student who are writing your graduation thesis. Continue to write the section from this thesis excerpt for approximately word_count words, maintaining formal academic structure and style. The original text:       . You are an IT student who are writing your graduation thesis. Extend the section by adding supporting detailed information for a thesis on       . The original text:       . You are a computer scientist who are very familiar with abstract writing. Add some concise concluding sentences to this partial abstract that highlights implications for future research. The original text:       . Continue the abstract by writing a closing statement that underscores the study’s contributions and potential applications. The original text:       .
AI-paraphrased Paraphrase the following thesis section in a clear academic tone, preserving citations and technical terms exactly. The original text:       . Reword this thesis excerpt to improve readability and maintain its scholarly voice, keeping all references unchanged. The original text:       . Rephrase this abstract in formal academic English, maintaining all original citations and technical accuracy. The original text:       . Paraphrase this abstract paragraph to enhance clarity and flow, ensuring all technical terms and citations remain intact. The original text:       .
Table 6: Samples of diverse prompt templates used to generate FAIDSet
Appendix C Analysis on similarity across models and model families
To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis using various perspectives: N-grams distribution, text length patterns, and semantic embedding visualization. These analyses allow us to examine similarities both within the same model family and across different model families, leading to a robust understanding of LLM "authorship" characteristics.
C.1 Text Distribution Between AI Families
We first examined the distribution of text length, measured in both word and character counts across outputs from five AI models: Llama-3.3-70B-Instruct-Turbo Dubey et al. (2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Team et al. (2024). Using 2000 arXiv prompt seeds, each model generated a corresponding output, and their length distributions were plotted.
From Figure 3, distinct patterns emerged between families. Gemini models consistently produced shorter and more compact outputs, while Llama and GPT models demonstrated more variance and a tendency toward longer completions. Despite using different versions (e.g., Gemini 2.0 vs. Gemini 1.5 Flash), the Gemini outputs remained closely grouped in terms of both word and character counts, suggesting a shared output generation strategy and stylistic consistency across the family. In contrast, Llama and GPT distributions showed greater separation and variability.
This reinforces the hypothesis that text length behavior is not only model-dependent but also family-coherent, with Gemini models forming a distinct cluster.
C.2 Text Distribution Between AI Models Within the Same Family
We performed N-grams frequency analysis on responses generated by three models within the Gemini family: Gemini 2.0, Gemini 2.0 Flash, and Gemini 1.5 Flash using 500 texts from the arXiv abstract dataset. Figure 5 highlights overlapping high-frequency tokens and similar patterns in word usage and phrase structure among the three models.
Despite minor differences in architectural speed (e.g., Flash vs. regular) or release chronology, the N-grams distributions show minimal divergence. Frequently used tokens, such as domain-specific terms and transitional phrases, appeared with nearly identical frequencies. This suggests that these models share similar decoding strategies and training biases, likely due to shared pretraining corpora and optimization techniques resulting in highly consistent stylistic patterns.
These intra-family similarities support treating model variants within a family as a unified authoring entity when performing analysis or authorship attribution.
C.3 Embedding Visualization and Semantic Cohesion
To delve into deeper semantic alignment, we visualized embeddings generated by unsupervised SimCSE XLM-RoBERTa-base model of generated texts from two model families: Gemini family (texts generated by Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash), and GPT-4o/4o-mini using PCA to reduce the high-dimensional embeddings into a space.
As shown in Figure 4, Gemini model embeddings form tight, overlapping clusters, indicating a high degree of semantic cohesion among their outputs. This clustering remains consistent across both sample sizes of 2000 texts. In contrast, GPT-4o/4o-mini embeddings occupy a separate region of the space, with greater dispersion and less overlap with Gemini clusters.
This visualization confirms that the Gemini family not only shares stylistic features but also maintains semantic coherence, distinguishing it from models of other families at a conceptual level.
C.4 Conclusion
The consistency observed across N-grams distributions, text length patterns, and semantic embeddings among Gemini models substantiates our decision to treat each AI family as an author. These models demonstrate coherent writing styles, shared lexical preferences, and tightly clustered semantic representations - hallmarks of unified authorship. Conversely, inter-family comparisons show clear separability, emphasizing the distinctiveness of each AI model family’s writing behavior.
(a) Word length distribution across five AI models.
(b) Character length distribution across five AI models.
Figure 3: Text length distributions in words and characters across Llama-3.3, GPT-4o/4o-mini, Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash.
(a) 2D embedding space.
(b) 3D embedding space.
Figure 4: Visualizations showing clustering behavior of Gemini model family (Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash) and GPT-4o/4o-mini using 2D and 3D embeddings with sample size of 2000 texts.
(a) Top 3-grams of Gemini 2.0 (500 samples).
(b) Top 3-grams of Gemini 2.0 Flash (500 samples).
(c) Top 3-grams of Gemini 1.5 Flash (500 samples).
Figure 5: Top 20 most common 3-gram from Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash using 500 sample prompts.
Appendix D Model selection for detector
Model #. Params Known Generators Unseen Generators
Acc $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$ Acc $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
RoBERTa-base 125M 80.09 76.22 0.7328 0.3778 73.45 69.10 0.8901 0.4320
FLAN-T5-base 248M 80.19 75.77 0.7783 0.3947 72.80 68.55 0.9123 0.4467
e5-base-v2 109M 81.53 77.90 0.8023 0.4086 74.21 70.15 0.8804 0.4392
Multilingual-e5-base 278M 91.41 90.82 0.3436 0.1732 85.32 84.50 0.5102 0.2543
XLM-RoBERTa-base 279M 91.90 90.63 0.2345 0.1190 86.75 85.20 0.4125 0.2104
Sup-SimCSE-RoBERTa-base 279M 81.22 78.88 0.7102 0.3619 74.00 71.30 0.8420 0.4251
UnSup-SimCSE-RoBERTa-base 279M 82.19 79.38 0.7156 0.3637 75.10 72.40 0.8305 0.4207
UnSup-SimCSE-XLM-RoBERTa-base 279M 92.12 91.75 0.1904 0.0958 87.45 86.90 0.3507 0.1802
Table 7: Model selection on known vs. unseen generators. The best results in each column are in bold.
To identify the best encoder for our classification task, we evaluated each candidate model on both known generators (FAIDSet testing data) and the new unseen-generator test set introduced in Section 3. Each transformer-based encoder was fine-tuned on FAIDSet training data and then used to predict labels on the two evaluation splits. Table 7 summarizes the accuracy, F1-macro, Mean Squared Error (MSE), and Mean Absolute Error (MAE) for each model under both conditions.
Base model comparison.
We first evaluated three popular monolingual models: RoBERTa-base, Flan-T5-base, and e5-base-v2 using the same training and evaluation splits. RoBERTa-base Liu et al. (2019b) and e5-base-v2 Wang et al. (2022) achieved the most balanced trade-off between classification accuracy and regression error (MSE, MAE), while Flan-T5-base Chung et al. (2022) lagged slightly in F1-macro. These results indicated that a stronger encoder backbone yields more robust performance and motivated exploration of multilingual variants for further gains.
Multilingual variants.
Next, we experimented with XLM-RoBERTa-base Conneau et al. (2019) and Multilingual-e5-base Wang et al. (2024a). Both models leverage cross-lingual pretraining, which in our scenario appears to improve representation of diverse linguistic patterns present in FAIDSet. In particular, XLM-RoBERTa-base delivered a substantial boost in all metrics, suggesting that its multilingual training enhances generalization even on primarily monolingual inputs.
Contrastive learning with SimCSE.
Finally, we integrated contrastive learning via SimCSE Gao et al. (2021) to further refine the sentence embeddings. We tested both supervised (trained on NLI data) and unsupervised (trained on Wikipedia corpus) SimCSE variants applied to RoBERTa-base. The unsupervised version outperformed its supervised counterpart, confirming prior findings that unsupervised SimCSE often yields stronger semantic encoders for downstream tasks. Encouraged by this, we adopted the unsupervised SimCSE approach on XLM-RoBERTa-base Vamvas and Sennrich (2023), which achieved the highest accuracy and lowest error rates across all configurations.
Based on these experiments, we selected the unsupervised SimCSE XLM-RoBERTa-base model for our final system.
Algorithm Known Generators Unseen Generators
Accuracy $\uparrow$ F1-macro $\uparrow$ Accuracy $\uparrow$ F1-macro $\uparrow$
k-Nearest Neighbors (KNN) 90.52 90.21 85.37 84.95
k-Means 88.13 87.48 80.22 79.81
Fuzzy k-Nearest Neighbors 95.18 95.05 93.31 93.25
Fuzzy C-Means 92.67 92.31 90.04 89.53
Table 8: Comparison of clustering algorithms on known vs. unseen generators. The best results are shown in bold.
Appendix E Ablation study
E.1 The need to use a vector database
We first used the trained model to classify the text and observed a significant performance decline from 90% when evaluating on unseen domains or text generated by novel models. This drop reflects the model’s limited ability to generalize beyond the distribution of its original training data. To address this, we propose integrating a vector database that stores dense embeddings of all collected examples, both training and incoming unlabeled data.
By indexing and retrieving semantically similar examples at inference time, the vector database serves as a flexible, scalable memory that bridges gaps between training and test distributions, enhancing the classifier’s resilience to domain shifts and unseen generators. Specifically:
  • Robust Domain Adaptation: New inputs are matched against a broad, continuously growing repository of embeddings, allowing the classifier to leverage analogous instances from related domains without full retraining.
  • Generator-Independent Coverage: As novel text generators emerge, their embeddings populate the database; retrieval naturally adapts to new styles or patterns by finding the closest existing vectors.
E.2 Clustering Algorithm Selection
To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies on the encoded vectors stored in our vector database. Each algorithm was tasked with grouping text samples into human-written, AI-generated, and human-AI collaborative categories, using both known-generator data (held-out from training) and entirely unseen generator data. Performance was measured in terms of accuracy and F1-macro score.
We encoded each text sample using our pretrained sequence-classification model’s penultimate layer, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three text-origin classes. From the results in Table 8, we draw some key findings:
  • Traditional algorithms show reasonable performance on held-out known generators but degrade notably on unseen generators.
  • Fuzzy C-Means leverages membership degrees to handle overlapping distributions, improving both accuracy and F1-macro by about 4% over k-Means, with smaller degradation on unseen data.
  • Fuzzy KNN combines local neighbor information with fuzzy membership, achieving the best overall performance: 95.2% accuracy and 95.1% F1-macro on known generators, and 93.3% accuracy and 93.3% F1-macro on unseen generators.
Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our full architecture.

### v2

@@
-mO Revanth[#1]
+FAID: Fine-Grained AI-Generated Text Detection Using Multi-Task Auxiliary and Multi-Level Contrastive Learning
 Abstract
-The growing collaboration between humans and AI models in generative tasks has introduced new challenges in distinguishing between human-written, AI-generated, and human-AI collaborative texts.
+The growing collaboration between humans and AI models in generative tasks has introduced new challenges in distinguishing between human-written, LLM-generated, and human–LLM collaborative texts.
 In this work, we collect a multilingual, multi-domain, multi-generator dataset FAIDSet.
-We further introduce a fine-grained detection framework FAID to classify text into these three categories, meanwhile identifying the underlying AI model family.
+We further introduce a fine-grained detection framework FAID to classify text into these three categories, and also to identify the underlying LLM family of the generator.
 Unlike existing binary classifiers, FAID is built to capture both authorship and model-specific characteristics.
 Our method combines multi-level contrastive learning with multi-task auxiliary classification to learn subtle stylistic cues.
-By modeling AI families as distinct stylistic entities, FAID offers improved interpretability.
-We incorporate an adaptation to address distributional shifts without retraining for unseen data.
-Experimental results demonstrate that FAID outperforms several baseline approaches, particularly enhancing the generalization accuracy on unseen domains and new AI models.
-It provide a potential solution for improving transparency and accountability in AI-assisted writing.
+By modeling LLM families as distinct stylistic entities, we incorporate an adaptation to address distributional shifts without retraining for unseen data.
+Our experimental results demonstrate that FAID outperforms several baselines, particularly enhancing the generalization accuracy on unseen domains and new LLMs, thus offering a potential solution for improving transparency and accountability in AI-assisted writing.
 1 Introduction
-AI has progressively evolved from an assistant tool to a creator or initiator, from helping polish papers and generate code to initiate proposals and draft essays, while humans increasingly serve as optimizers and reviewers.
-In such deeply collaborative human-AI settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al.
+LLMs have evolved from an assistant tool to a creator or initiator, from helping polish papers and generate code to initiate proposals and draft essays, while humans increasingly serve as optimizers and reviewers.
+In such deeply collaborative human–LLM settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al.
 (2025).
-This work aims to identify human involvement by distinguishing the origin of a given text: fully AI-generated, fully human-written, or collaboratively produced, in multilingual context.
-Numerous studies have explored multilingual AI-generated text detection, but most focus either on binary detection: human vs.
-AI Wang et al.
+This work aims to identify human involvement by distinguishing the origin of a given text (in a multilingual context): (a) fully LLM-generated, (b) fully human-written, or (c) collaboratively produced.
+Numerous studies have explored multilingual LLM-generated text detection, but most have focused either on binary detection, i.e., human vs.
+LLM Wang et al.
 (2024d, b), or on fine-grained detection limited to English Abassy et al.
@@
 (2024a).
-Moreover, both struggle with generalization to unseen domains, languages, and new AI generators Wang et al.
+Moreover, both struggle with generalization to unseen domains, languages, and LLM generators Wang et al.
 (2024c); Li et al.
 (2024).
-Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain and multi-generator dataset for fine-grained detection, i.e., identifying a text into three categories: AI-generated, human-written, and human-AI collaborative; and (ii) introducing a detection framework FAID to ameliorate poor generalization performance.
-[^1] Our code and data are accessible at github.com/ngocminhta/FAID.
-Our dataset focuses on academic field including domains of research paper abstracts, student theses, and scholarly reports, and involve recent LLM families, e.g., GPT, Gemini, DeepSeek and Llama OpenAI (2024); Team et al.
+Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain and multi-generator dataset, FAIDSet, for fine-grained detection, i.e., identifying a text into three categories: LLM-generated, human-written, and human–LLM collaborative; and (ii) introducing a detection framework FAID to improve generalization performance.
+Our dataset FAIDSet focuses on the academic field including paper abstracts, student theses, and scholarly reports, and contains generation by a variety of families of LLMs, e.g., GPT, Gemini, DeepSeek, and Llama OpenAI (2024); Team et al.
 (2024); DeepSeek-AI (2024); Dubey et al.
 (2024).
-FAID treats each AI model as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features.
-This was achieved by optimizing a language encoder with multi-level author relationship (e.g., stylistic similarity between texts from the same LLM family is greater than the similarity between a human and an AI) to capture author-specific distinguishable signals of an input text based on contrastive learning, along with the task of fine-tuning a classifier to recognize its origin.
-This multitask learning process forces the encoder to reorganize the hidden space, representations of texts from the same authors distributed closer than from different authors.
-Practically, given that generations produced by models developed by the same company tend to have similar writing styles due to similar model architecture, training data and training strategy (more details in Appendix C), we consider each AI family as an author.
-This can also help the detector acquire prior knowledge of future LLMs from the same company.
-Experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations.
+Our detection framework FAID treats each LLM as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features.
+We achieve this by optimizing a language encoder with multi-level author relationship (e.g., the stylistic similarity between texts from the same LLM family is greater than the that between a human and an LLM) to capture author-specific distinguishable signals of an input text using contrastive learning, along with the task of fine-tuning a classifier to recognize the input text’s origin.
+This multitask learning process forces the encoder to reorganize the hidden space, so that the representations of texts by the same authors are distributed closer than such by different authors.
+In practice, given that generations produced by LLMs from same company tend to have similar writing styles due to similar model architecture, training data, and training strategy (see Appendix D for more detail), we consider each LLM family as an “author.” This can also help the detector acquire prior knowledge of future LLMs from the same company.
+Our experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations.
 Our contributions can be summarized as follows:
-  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained AI-generated text detection with 84k examples.
-  • We propose a detection framework FAID to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in hidden representation.
-  • Experimental results show that FAID outperforms baseline detectors, particularly over unseen domains and generators.
-Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to others with ground truth authorship in the database, implicitly enhancing interpretability with cues explaining why a text is assigned to a specific class.
-2 Backgrounds
-Advanced LLMs have changed how we create and refine text.
-Instead of originating every word, humans now more often act as post-editors or reviewers, stepping in after AI drafts the initial content.
-This collaborative workflow spans fields from academic publishing and journalism to education and social media, making hybrid human-AI texts the new norm Cheng et al.
+  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained LLM-generated text detection with 84k examples.
+  • We propose a detection framework FAID to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in the hidden representation.
+  • We show that FAID outperforms baseline detectors, particularly over unseen domains and generators.
+Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to other texts from our database, labeled with ground truth authorship labels.
+2 Background and Related Work
+Advancements in LLMs have fundamentally reshaped the process of text production and refinement.
+Rather than originating every word independently, people increasingly assume the roles of post-editors or reviewers, intervening after an initial draft has been generated by an LLM.
+This collaborative paradigm extends across diverse domains, including academic publishing, journalism, education, and social media, thus making hybrid human–AI authorship an emerging norm Cheng et al.
 (2024); Lee et al.
@@
 AI.
-Texts can be fully human-written, fully AI-generated, or collaborative text (e.g., human-written, AI-polished; human-initiated, AI-continued; AI-generated, human-edited; and deeply mixed text) Artemova et al.
+Texts can be fully human-written, fully AI-generated, or collaborative text (e.g., human-written, AI-polished; AI-generated, human-edited; and deeply mixedArtemova et al.
 (2025).
-This demands the fine-granularity authorship detection, and even tracing back to a specific model family, to measure human contributions Hutson (2025).
-Then, clear attribution of AI involvement helps uphold honesty in research and preserve scientific integrity.
-Responding to these challenges, we collect a new dataset, and propose a detection framework which generalizes well to new domains, languages, and models, showing high accuracy and reliability.
+Addressing these concerns necessitate fine-granularity authorship detection, and even tracing back to a specific LLM family, in order to assess the extent of human contribution Hutson (2025).
+Ensuring transparent disclosure of LLM involvement is critical for upholding research honesty and preserving the integrity of research.
+In response to these challenges, we collect a new dataset, FAIDSet, and propose a detection framework, FAID, which generalizes well to new domains, languages, and models, achieving consistently high accuracy and reliability.
 2.1 Fine-Grained AI-generated Text Datasets
-Many prior studies have explored fine-grained AI-generated text datasets across various forms of human-AI collaboration as shown in Table 3, for example, MixSet (Zhang et al., 2024b), and Beemo (Artemova et al., 2025).
-More datasets and detailed information regarding the label space, and the coverage of domains, languages and LLMs for each are summarized in Table 3.
-Despite these contributions, they have been limited to English.
-There is a noticeable gap in the availability of large-scale, fine-grained multilingual AI-generated text detection datasets.
-To bridge this gap, we collect a multilingual fine-grained AI-generated text detection dataset.
-It encompasses 84k texts generated by the latest LLMs and includes diverse forms of human-AI collaborative generations.
+Many prior studies have explored fine-grained AI-generated text datasets across various forms of human–AI collaboration, e.g., MixSet (Zhang et al., 2024b), and Beemo (Artemova et al., 2025).
+See Table 3 in the Appendix, for a discussion of more datasets and detailed information regarding the label space, and the coverage of domains, languages and LLMs for each one.
+However, all these studies were limited to English.
+A substantial gap remains in the availability of large-scale, fine-grained multilingual LLM-generated text detection datasets.
+To bridge this gap, we collect a multilingual fine-grained LLM-generated text detection dataset, which encompasses 84k texts generated by the latest LLMs and includes diverse forms of human–LLM collaborative generations.
 This dataset can facilitate the development of more robust and generalizable detection models that are capable of handling complex multilingual collaborative scenarios.
 2.2 Generalization of AI-Generated Text Detection
-Recent studies such as M4GT Wang et al.
-(2024c) and LLM-DetectAIve Abassy et al.
-(2024) have highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators.
-Many detection methods show a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors to deal with real-world scenarios and evolving LLMs outputs.
+A recent study, M4GT-Bench, Wang et al.
+(2024c) has highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators.
+Many detection methods have shown a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors to deal with real-world scenarios and evolving LLMs outputs.
 Various approaches have been proposed to improve generalization.
@@
 LLM-DetectAIve Abassy et al.
-(2024) adopts fine-grained multi-way classification and incorporates domain-adversarial training to reduce overfitting, yet its generalization to truly unseen domains and generators remains limited.
-Also, its current version lacks multilingual support.
+(2024) adopts fine-grained classification and incorporates domain-adversarial training to reduce overfitting; yet, its generalization to unseen domains and generators remains limited; moreover, its current version lacks multilingual support.
 SeqXGPT Wang et al.
-(2023) focused on sentence-level detection by combining log-probability features with convolutional and self-attention mechanisms, which helps capture subtle mixed-content signals and improves generalization across input styles.
+(2023) focuses on sentence-level detection by combining log-probabilities with convolutional and self-attention mechanisms, which helps to capture subtle mixed-content signals and improves generalization across input styles.
 However, its reliance on specific model features and limited semantic representation constrains adaptability to new generators and domains.
-A recent representative work, the DeTeCtive framework Guo et al.
-(2024), introduces multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, though it primarily focuses on binary classification and does not fully address human-AI collaborative texts.
+Finally, the DeTeCtive framework Guo et al.
+(2024), introduces multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, but it primarily focuses on binary classification and does not fully address human–AI collaborative texts.
 2.3 Contrastive Learning for AI-Generated Text Detection
-Contrastive learning has been extensively used to improve the sentence representation by pulling semantically similar sentences closer and pushing dissimilar ones apart, reorganizing the hidden space of sentence representation by semantics.
-For example, a representative work SimCSE regarded an input sentence under dropout noise as its semantically similar counterpart (i.e., positive pairs), and trained the encoder to close their distance.
-Additionally, they natural language inference (NLI) pairs, treating entailment pairs as positives and contradiction pairs as hard negatives (Gao et al., 2021).
-For negative pairs with contradictory meanings, the embeddings were trained to be distant in the latent space.
-Similarly, DeCLUTR Giorgi et al.
-(2021) constructed positive pairs by extracting different spans from the same texts, and negative pairs are sampled from different texts.
-We adopt this philosophy to reorganize the latent space based on writing styles, pulling human-written texts to cluster together while remaining distant from AI-generated ones.
-Analogous to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement in collaborative texts, ranging from 0 (fully AI-generated) to 1 (fully human-written).
-DeTeCtive also took advantage of contrastive learning and was used in binary AI-generated text detection Guo et al.
-(2024).
-Based on a multi-task framework, DeTeCtive was trained to learn writing-style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs.
+Contrastive learning has been extensively used to improve sentence representation by pulling semantically similar sentences closer and pushing dissimilar ones apart.
+SimCSE treats a sentence subjected to dropout noise as a semantically similar counterpart (i.e., a positive pair) and trains the encoder to minimize the distance between the original and the noise sentence.
+It further leverages natural language inference pairs, considering entailment pairs as positives and contradiction pairs as hard negatives (Gao et al., 2021).
+It is then trained to maximize the separation between negative pairs in the embedding space.
+DeCLUTR Giorgi et al.
+(2021) constructed positive pairs by extracting different spans from the same texts, and by sampling negative pairs from different texts.
+We adopt the same core idea, but reorganize the latent space based on writing style, pulling human-written texts to cluster together while remaining distant from LLM-generated ones.
+Similarly to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement in collaborative texts, ranging from 0 (fully LLM-generated) to 1 (fully human-written).
+Another work, DeTeCtive Guo et al.
+(2024), also took advantage of contrastive learning and was used in binary AI-generated text detection.
+Based on a multi-task framework, it was trained to learn writing-style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs.
 AI) to capture distinguishable signals.
-For inference, the pipeline encoded the input text as a hidden vector and applied dense retrieval to match the cluster, depending on the stylistic similarity with previously indexed training feature database.
+For inference, their pipeline encoded the input text as a hidden vector and applied dense retrieval to match the cluster, depending on the stylistic similarity with respect to a database of previously indexed training features.
 Additionally, instead of retraining the model when faced with new data, they encoded new data with the trained encoder to obtain embeddings, and then added them to the feature database to augment.
-This largely improved the generalizability in unseen domains and new generators.
-However, this approach only considers two types of text: human-written and AI-generated, without further consideration of human-AI collaborative texts.
-Our work fills this gap with the goal of improving generalization performance for fine-grained AI-generated text detection.
+This largely improved the generalizability to unseen domains and new generators.
+However, this approach distinguishes only between two categories of text (human-written vs.
+LLM-generated) while overlooking the increasingly prevalent class of human–LLM collaborative texts.
+Our work bridges this gap with the aim of enhancing generalization performance for fine-grained LLM-generated text detection.
 3 FAIDSet
-To fill the gap of fewer publicly-available multilingual fine-grained AI-generated text datasets, we collected a multilingual, multi-domain and multi-generator AI-generated text dataset FAIDSet.
-It contains texts generated by LLMs, written by humans and collaborated by both humans and LLMs, resulting in a total of $\sim$84k examples.
-See Appendix B for more details about FAIDSet.
-FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages including Vietnamese and English.
+We collected a new multilingual, multi-domain and multi-generator LLM-generated text dataset, FAIDSet.
+It contains texts generated by LLMs, written by humans and collaborated by both humans and LLMs, resulting in a total of 84k examples; see Appendix B for more details about FAIDSet.
+FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages (Vietnamese and English).
 We collected student theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^1] https://vjol.info.vn/ (VJOL).
 Models and Label Space.
-We used the following multilingual LLMs families to produce both AI and human-AI collaborative texts: GPT-4/4o, Llama-3.x, Gemini 2.x, and Deepseek V3/R1.
-Regarding the human-AI collaborative text, we include AI-polished, AI-continued, and AI-paraphrased where the models are requested to polish or paraphrase inputs while ensuring the accuracy of any figures and statistics.
+We used the following multilingual LLM families to produce LLM and human–LLM collaborative texts: GPT-4/4o, Llama-3.x, Gemini 2.x, and Deepseek V3/R1.
+Regarding the human–LLM collaborative text, we include LLM-polished, LLM-continued, and LLM-paraphrased, where the models are requested to polish or paraphrase inputs while ensuring the accuracy of any figures and statistics.
 Diverse Prompt Strategies.
-We generate data with diverse tones and contexts while ensuring the accuracy of the content.
-Depending on the data source and context, we craft prompts to create varied outputs suitable for different real-world application scenarios.
-We generate responses with different tones using prompts such as "You are an IT student…" and "…who are very familiar with abstract writing…", which are described more in Appendix B.2.
+We generated data with diverse tones and contexts while ensuring the accuracy of the content.
+Depending on the data source and context, we crafted prompts to create varied outputs suitable for different real-world scenarios.
+We generated responses with different tones using prompts such as “You are an IT student…” and “…who are very familiar with abstract writing…”.
+See Appendix B.2 for the full list.
+While FAIDSet does not capture the full natural diversity of in-the-wild LLMs outputs, the controlled setup enables us to systematically model stylistic and collaborative signals across multiple languages and families.
+This makes FAIDSet both a reproducible training resource and a benchmarking corpus: it provides reliable supervision for developing detectors while also serving as a testbed for assessing generalization to more diverse, unseen scenarios.
 Quality Control.
-Given that the task of AI-generated text detection naturally has gold labels when collecting data, we perform quality control by randomly sampling 10-20 instances for each domain, source, and generative model in data collection, and manually assess whether or not the generated texts are fluent and logically reasonable, to avoid incoherence, disorganization, or illogical texts.
-Such cases introduce remarkably distinguishable signals from human-written text, affecting the data quality.
-We adjust either the prompts or the generation parameters to address these problems.
+To avoid bad machine-generated texts, which can introduce remarkably distinguishable signals, we performed quality control by randomly sampling 10-20 instances for each domain, source, and LLM generator.
+Manual inspection focused on fluency, coherence and factual plausibility.
+Overall, the generated texts demonstrated high linguistic quality with most outputs being fluent and logically reasonable.
+Nevertheless, occasional issues such as repetitive phrasing, incomplete reasoning or overuse of formal expressions were observed.
+In that cases, we adjusted the prompts (e.g., by specifying desired length or style) or refined generation parameters to improve diversity and logical consistency.
+After these adjustments, the quality across generators and domains was found to be stable and satisfactory.
 4 Methodology
-In this section, we first clarify task definition and then elaborate our detection method.
-4.1 Task Definition
-In this work, we formulate the task as a three-class classification problem: human-written, AI-generated, and human-AI collaborative texts.
-The human-AI collaborative category involves a range of interaction between humans and AI systems, such as human-written, AI-polished; human-initiated, AI-continued; human-written, AI-paraphrased; AI-generated, AI-humanized, etc.
-Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, we do not exhaust all patterns.
-Instead, we consolidate all forms of human-AI collaboration into a single label to maintain practical simplicity and model generalizability.
-This approach reflects real-world usage, where using AI tools to enhance clarity or expression is increasingly common and often ethically acceptable.
-We also investigate the dataset and find that the AI models within the same family tend to have similar writing styles and text distributions, due to their shared training data and architecture (see Appendix C).
-Therefore, we consider each model’s family to be an author with unique writing styles.
+Task Definition
+Our task is a three-class classification problem: human-written vs. LLM-generated vs. human–LLM collaborative text detection.
+The human–LLM collaborative category involves a range of interaction between humans and LLM systems, such as human-written, LLM-polished; human-initiated, LLM-continued; human-written, LLM-paraphrased, etc.
+Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, this is not exhaustive.
+Instead, we consolidate all forms of human–LLM collaboration into a single label to maintain practical simplicity and model generalizability.
+This reflects real-world usage, where using LLM tools to enhance clarity or expression is increasingly common and often ethically acceptable.
+Our analysis of the dataset revealed that the LLM models within the same family tend to have similar writing style and text distributions, due to their shared training data and architecture (see Appendix D).
+Thus, we consider each model’s family to be an “author” with a unique writing style.
 4.1 Framework Overview
 Figure 1: Training architecture.
-Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix D) based on the human, human-AI and AI texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther.
+Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix E) based on the human, human–LLM and LLM-generated texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther.
 We train an encoder that can represent text with distinguishable signals to discern authorship of text.
-We aim to enable the encoder to capture multi-level similarities between authors (AI, human and human-AI) as Eq (6), forcing model to learn features that can differentiate distinct authors in a high-dimensional hidden vector space.
-We denote $\mathrm{S}_{c}$ as cosine similarity, $\phi(\cdot)$ is the encoder function and $P_{i}$, $P_{j}$, ($51\leq i\leq j<$ 5) are distributions of different text subsets.
-We have:
+Here, we shift the focus to the detector, an encoder-based model that forms the core of FAID.
+The detector encodes each text into a high-dimensional embedding space to quantify cross-source similarity, enabling the study of both intra- and inter-family relationships.
+It is trained to capture multi-level similarities between authors by learning a representation space where closely related sources (e.g., LLMs from the same family) form tighter clusters, while dissimilar ones (e.g., human vs.
+LLM) are pushed farther apart.
+Let $\mathrm{S}_{c}$ be cosine similarity, $\phi(\cdot)$ be the encoder function, and $P_{i}$, $P_{j}$, ($1\leq i\leq j<5$) be the distributions of different text sources.
+We aim for the model to encode representations that satisfy the following constraint:
 $\begin{split}\mathbb{E}_{x\in P_{i},y\in P_{j}}&\left[\mathrm{S}_{c}\left(\phi(x),\phi(y)\right)\right]\\ &\geq\mathbb{E}_{x\in P_{i},y\in P_{j+1}}\left[\mathrm{S}_{c}\left(\phi(x),\phi(y)\right)\right]\end{split}$ (1)
-where $P_{1}$ corresponds to the distribution generated by a particular LLM family, $P_{2}$ to the distribution generated by any LLMs, $P_{3}$ to the distribution of a collaborative text generated by humans and the LLM family of $P_{1}$, $P_{4}$ to the distribution of collaborative text generated by humans and any LLM families, and $P_{5}$ to the distribution of human-written text.
-To clarify the rationale for configuring FAID to expect that the similarity of a text $x$ (from lower-level distributions $P_{1}$ or $P_{2}$) with samples from $P_{3}$ is generally greater than or equal to its similarity with samples from $P_{4}$, consider the following.
+where $P_{1}$ corresponds to the distribution generated by a particular LLM family, $P_{2}$ is the distribution generated by any LLM, $P_{3}$ is the distribution of collaborative text generated by human and a LLM family of $P_{1}$, $P_{4}$ is the distribution of collaborative text generated by humans and any LLM families, and $P_{5}$ is the distribution of human-written text.
+To clarify the rationale for configuring FAID to expect that the similarity of a text $x$ (from lower-level distributions $P_{1}$ or $P_{2}$) with samples from $P_{3}$ is generally greater than or equal to its similarity with samples from $P_{4}$, consider the following:
   • If $x\in P_{1}$ ($x$ is generated by a particular LLM): Let $y_{\textit{LHS}}$ be drawn from $P_{3}$ and $y_{\textit{RHS}}$ be drawn from $P_{4}$.
-Naturally, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ should be greater than $\mathrm{S}_{c}(x,y_{\textit{RHS}})$.
+Naturally, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is greater than $\mathrm{S}_{c}(x,y_{\textit{RHS}})$.
 This is because $P_{3}$ contains texts that share a direct LLM family origin with $x$.
-  • If $x\in P_{2}$ ($x$ is generated by any LLM): In this scenario, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$.
-Since $x$ can originate from any LLM, it does not inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
-This configuration aims to ensure that closeness in distribution corresponds to heightened similarity after encoding, encouraging the model to discern fine-grained multi-level relations.
+  • If $x\in P_{2}$ ($x$ is generated by any LLM): Here, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$.
+Since $x$ can originate from any LLM, it doesn’t inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
+This configuration aims to ensure that closeness in distribution corresponds to higher similarity after encoding, thus encouraging the model to discern fine-grained multi-level relations.
 4.2 Multi-level Contrastive Learning
-Given a dataset with $N$ samples, the $i^{th}$ sample is denoted as $T_{i}$ and assigned with three-level labels indicating its source.
-  • $x$: if $T_{i}$ is fully AI-generated, $x_{i}=0$, otherwise $x_{i}=1$;
-  • $y$: if $T_{i}$ is fully human-written, $y_{i}=0$, otherwise $y_{i}=1$;
-  • $z$: an indicator of a specific LLM family.
+Given a dataset with $N$ samples, each sample is a text unit (paragraph/segment). The $i^{th}$ sample is denoted as $T_{i}$. Per sample, we assigned three-level labels indicating its source:
+  • $x_{i}\in\{0,1\}$: if $T_{i}$ is fully LLM-generated, $x_{i}=0$, otherwise $x_{i}=1$;
+  • $y_{i}\in\{0,1\}$: if $T_{i}$ is fully human-written, $y_{i}=0$, otherwise $y_{i}=1$;
+  • $z_{i}$: an indicator of a specific LLM family.
 The encoder $\phi(\cdot)$ represents the sample $T_{i}$ in a $d$-dimensional vector space $\mathbb{R}^{d}$.
 Then we calculate the cosine similarity between two samples $T_{i}$ and $T_{j}$, denoted by: $\sigma(i,j)=\mathrm{S}_{c}\left(\phi(T_{i}),\phi(T_{j})\right)$.
-For AI-generated text ($x=0$), the similarity between $T_{i}$ and another AI-generated text $T_{j}$ is greater than that with a human-written or collaborative text $T_{k}$ ($x=1$):
+For LLM-generated text, the similarity between $T_{i}$ and another LLM-generated text $T_{j}$ is greater than that with a human-written or collaborative text $T_{k}$:
 $\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,x_{i}=x_{j},x_{k}=1$ (2)
-If $x=0$, then the text is fully AI-generated.
-For this case, we do not consider $y$ since AI-generated text is considered as non-AI-collaboration.
-We can imply that the similarity between two texts written by the same AI is higher than that of two different AI families.
+If $x_{i}=0$, then the text sample is fully LLM-generated.
+For this case, we do not consider $y_{i}$ since LLM-generated text is considered as non-LLM-collaboration.
+We can imply that the similarity between two texts written by the same LLM is higher than that of two LLM families.
 Hence:
 $\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,z_{i}=z_{j},z_{i}\neq z_{k}$ (3)
-The reverse condition (with $y=1$) is also true. We can conclude that:
+The reverse condition is also true. We can conclude that:
 $\sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i}=y_{j},y_{i}\neq y_{k}$ (4)
-For cases where all samples are human-AI collaborative texts ($x,y=1$), two texts with involvements of the same AI family tend to be more similar than the ones that involve contributions from different families.
-That is:
+For cases where all samples are human–LLM collaborative, two texts created by the same LLM family tend to be more similar than the ones that involve contributions from different families. That is:
 $\sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i,j,k}=1,z_{i}=z_{j}\neq z_{k}$ (5)
 Combining all, the text representation is learned with the following constraints:
-$\begin{cases}\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,x_{i}=x_{j}\neq x_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=0,z_{i}=z_{j}\neq z_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=1,x_{i}=x_{j}\neq x_{k};\\ \sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i}=y_{j}\neq y_{k};\\ \begin{split}\sigma(i,j)>\sigma(i,k),\forall&x_{i}=1,y_{i,j,k}=1,\\ &z_{i}=z_{j},z_{i}\neq z_{k}\\ \end{split}\end{cases}$
+[…]
 To enforce the similarity constraints outlined in Eq (6), we build upon the SimCLR framework Chen et al.
@@
 Departing from traditional contrastive losses that rely on a single positive sample, our approach considers a group of positive instances that satisfy specific criteria.
-The similarity between the anchor and positive samples is computed as the average similarity across this entire positive set.
+The similarity between the anchor and the positive samples is computed as the average similarity across this entire positive set.
 For negative samples, we follow the methodology used in SimCLR.
@@
 $\begin{split}\mathcal{L}_{mcl}=&\sum\limits_{i=1}^{N}\left[x_{i}\left(\alpha\mathcal{L}_{q_{i},1}+\beta\mathcal{L}_{q_{i},2}\right)+\right.\\ &\left.\left(1-x_{i}\right)\left(\gamma\mathcal{L}_{q_{i},3}+\delta\mathcal{L}_{q_{i},4}+\zeta\mathcal{L}_{q_{i},5}\right)\right]\end{split}$ (8)
-Therefore, due to the last inequality in Eq (6) only specifying a case ($y=1$), and the other cases considering both values for $y$, so we have $\zeta=2\gamma=2\delta$.
+Therefore, due to the last inequality in Eq (6) only specifying a case ($y_{i,j,k}=1$), and the other cases considering both values for $y$, so we have $\zeta=2\gamma=2\delta$.
 Also, to maintain the equilibrium, we need to keep $\alpha+\beta=\gamma+\delta+\zeta$.
 We set $\gamma=\delta=1$, then $\zeta=2,\alpha=\beta=2$.
-This approach encourages the model to capture subtle and detailed features from different sources.
-As a result, the model becomes more adept at recognizing variations in writing styles.
-This capability enhances its accuracy and strengthens its ability to generalize when detecting AI-generated text.
+This encourages the model to capture subtle and detailed features from different sources.
+As a result, it becomes more adept at recognizing variations in writing styles.
+This capability enhances its accuracy and strengthens its ability to generalize when detecting LLM-generated text.
 4.3 Multi-Task Auxiliary Learning
@@
 This joint learning process helps the model develop more general and distinctive features.
-As a result, it improves the model’s ability to generalize to new data.
+Therefore, it improves the model’s generalizability to new data.
 Building on the previously described contrastive learning framework, we extend the encoder by attaching an MLP classifier at its output layer.
-This classifier is responsible for binary classification, determining whether a given query text was written by a human or generated by a LLM.
-Denote that the probability of $i^{th}$ sample with label $x_{i}=0$ is $p_{i}$.
-To train this component, we apply a cross-entropy loss function, denoted as $\mathcal{L}_{ce}$, defined as follows:
+This classifier is responsible for binary classification, determining whether a given text was written by a human or a LLM.
+Let the probability of $i^{th}$ sample with label $x_{i}=0$ be $p_{i}$.
+To train this component, we apply a cross-entropy loss function, denoted as $\mathcal{L}_{ce}$, and defined as follows:
 $\mathcal{L}_{ce}=-\dfrac{1}{N}\sum\limits_{i=1}^{N}x_{i}\log p_{i}+\left(1-x_{i}\right)\log\left(1-p_{i}\right)$ (9)
@@
 $\mathcal{L}=\mathcal{L}_{ce}+\mathcal{L}_{mcl}$ (10)
-Figure 2: Inference architecture.
-(a) embed the input text into embedding vector using the fine-tuned encoder; (b) use Fuzzy KNN algorithm to cluster, retrieving which cluster the input text belongs to (see more in Appendix E).
-(c) The stored vector database $\mathcal{VD}$ was created by saving all embeddings of texts in training and validation sets using the fine-tuned encoder.
-If the input text is unseen, we embed the unseen text and save into a temporary vector database $\mathcal{VD}^{\prime}$, enhancing the generalization of the detector.
+Figure 2: Inference architecture: (a) embed the input text into embedding vector using the fine-tuned encoder, (b) use Fuzzy kNN to cluster, retrieving which cluster the input text belongs to (see more in Appendix F), (c) the stored vector database $\mathcal{VD}$ was created by saving all embeddings of texts in training and validation sets using the fine-tuned encoder.
+If the input text is unseen, we embed it and save it into a temporary vector database $\mathcal{VD}^{\prime}$, enhancing the generalization of the detector.
 4.4 Handling Unseen Data without Retraining
-Unseen data, whether from an unseen domain or generated by an unfamiliar model, remains a significant challenge even for state-of-the-art AI-generated text detection methods, due to the increasing capabilities of large language models.
-We tried to use the model classifier on its own, but we ended up by using a vector database along with Fuzzy k-Nearest Neighbors, which is illustrated in Figure 2.
-The results were written in Appendix E.
-Specifically, when dealing with unseen data, we use our trained model to embed these samples and add them to our existing vector database.
+Unseen data, whether from an unseen domain or an unfamiliar generator, remains a significant challenge even for state-of-the-art LLM-generated text detection methods, due to the increasing capabilities of LLMs.
+We tried to use the model classifier on its own, but we ended up using a vector database along with Fuzzy k-Nearest Neighbors, as illustrated in Figure 2.
+The results were written in Appendix F.
+Specifically, when dealing with unseen data, we use our model to embed these samples and add them to our existing vector database.
 Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without the need retraining.
 5 Experiments
-In this section, we first describe datasets and baselines used, followed by two experiments evaluating FAID in classifying a text between human, AI and human-AI; and in identifying specific generators.
+In this section, we describe the datasets and baselines we used, followed by two experiments evaluating FAID: (i) classify a text as human, LLM, and human–LLM, and (ii) identify specific generators.
 Scenario Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
-In-domain, Known generator FAIDSet LLM-DetectAIve 94.34 94.45 93.79 94.10 0.1888 0.1107
-SVM (N-grams) 84.41 84.53 82.93 83.62 0.5630 0.2916
+In-domain,
+Known generator
+FAIDSet LLM-DetectAIve 94.34 94.45 93.79 94.10 0.1888 0.1107
 T5-Sentinel 93.31 94.92 93.10 93.15 0.2104 0.1101
+SeqXGPT 85.77 85.49 86.02 84.69 0.5593 0.2844
 FAID 95.58 95.78 95.33 95.54 0.1719 0.0875
 LLM-DetectAIve LLM-DetectAIve 95.71 95.78 95.72 95.71 0.1606 0.1314
-SVM (N-grams) 83.43 76.09 71.84 73.60 0.2878 0.2064
 T5-Sentinel 94.77 94.70 92.60 93.60 0.1663 0.1503
+SeqXGPT 81.48 78.72 74.91 76.71 0.3141 0.2255
 FAID 96.99 95.29 88.14 91.58 0.1561 0.0754
 HART LLM-DetectAIve 94.39 94.25 94.33 94.29 0.3244 0.1789
-SVM (N-grams) 58.93 59.75 61.07 60.30 1.2334 0.6849
 T5-Sentinel 86.68 87.25 87.69 87.38 0.4339 0.2334
+SeqXGPT 63.12 64.01 65.27 64.05 1.0057 0.5982
 FAID 96.73 97.61 98.05 97.80 0.4631 0.1806
-Unseen domain, Unseen generator Unseen domain LLM-DetectAIve 52.83 47.31 64.62 53.28 0.4733 0.4722
-SVM (N-grams) 39.28 43.13 29.41 22.77 0.7689 0.6611
+Unseen domain,
+Unseen generator
+Unseen domain LLM-DetectAIve 52.83 47.31 64.62 53.28 0.4733 0.4722
 T5-Sentinel 55.56 49.54 66.67 55.34 0.4444 0.4444
+SeqXGPT 40.60 43.81 31.87 36.72 0.8021 0.7028
 FAID 62.78 70.73 71.77 69.46 0.4514 0.4486
-Unseen generators LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
-SVM (N-grams) 74.19 61.88 50.17 55.17 0.4324 0.3162
+Unseen
+generators
+LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
 T5-Sentinel 85.95 85.77 84.59 85.16 0.3648 0.2419
+SeqXGPT 72.04 60.33 48.94 54.12 0.4590 0.3380
 FAID 93.31 92.40 94.44 93.25 0.1691 0.1167
-Unseen domain and Unseen generators LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
-SVM (N-grams) 34.50 39.38 26.67 21.70 0.9186 0.7429
+Unseen domain
+and Unseen
+generators
+LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
 T5-Sentinel 57.07 49.82 66.61 55.45 0.4314 0.4300
+SeqXGPT 40.71 47.95 35.21 40.09 0.8753 0.7086
 FAID 66.55 74.44 73.57 72.58 0.3939 0.3167
 Table 1: Performance with three labels.
-For in-domain and known generator datasets, we train detectors on training set and evaluate using test set.
-For unseen data evaluation, we use the detector trained on FAIDSet, and evaluate on the unseen datasets.
+For in-domain and known generator datasets, we train the detectors on the training set and evaluate them on the test set.
+For unseen data evaluation, we use the detector trained on FAIDSet and evaluate on the unseen datasets.
 The best results are in bold and the second best are underlined.
 5.1 Datasets
-In addition to FAIDSet, for in-domain evaluation, we used another two datasets:
+In addition to FAIDSet, for in-domain evaluation, we used two additional datasets:
 LLM-DetectAIve Abassy et al.
-(2024) encompasses various domains including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews.
-The original human-written and machine-generated texts were augmented using a variety of LLMs (e.g., Llama 3, Mixtral, Gemma) to create 236,000-example dataset with two new labels: machine-written then machine-humanized, and human-written then machine-polished.
+(2024) encompasses various domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews.
+The original human-written and machine-generated texts were augmented by multiple LLMs to create a 236k-example dataset with two new labels: (imachine-written then machine-humanized, and (iihuman-written then machine-polished.
 HART Bao et al.
-(2025) is a dataset spans domains of student essays, arXiv abstracts, story writing, and news articles.
-It has four categories: human-written texts, LLMs-refined texts, AI-generated texts, and humanized AI-generated texts.
-They further generated 21,500 cases based on this dataset for unbalanced labels.
-To evaluate generalization ability of FAID on unseen scenarios, we collected the following data:
-Unseen domain: We create a dataset consisting of 150 IELTS Writing essays from Kaggle ibrahimmazlum (2023), where all original texts are human-written.
-These essays were then used to generate human-AI collaborative and AI-generated texts, using the same models with FAIDSet.
-Unseen generators: We selected 150 human-written abstracts from the FAIDSet test set and generated data for remaining labels using the three new LLM families: Qwen 2, Mistral, and Gemma 2.
-Unseen domain & generators: Based on the human-written IELTS essays above, instead of the original four LLMs, we used the same LLM families as Unseen generator test set to generate data for the AI and human-AI labels, following the same process as with the unseen domains.
+(2025) with 21,500 examples including student essays, arXiv abstracts, story writing, and news articles.
+It contains four categories: human-written, AI-refined, AI-generated, and humanized AI-generated texts.
+The authors further expanded the dataset by creating additional instances with unbalanced label distributions.
+To evaluate the generalizability of FAID on unseen scenarios, we collected the following data:
+Unseen domain: We created a dataset consisting of 150 IELTS Writing essays from Kaggle ibrahimmazlum (2023), where all original texts are human-written.
+We used these essays to generate human–LLM collaborative and LLM-generated texts with the same models with FAIDSet.
+Unseen generators: We selected 150 human-written abstracts from the FAIDSet test set and generated data for the remaining labels using three new LLM families: Qwen, Mistral, and Gemma.
+Unseen domain & generators: Based on the human-written IELTS essays above, we used the same LLM families as for the unseen generator test set to generate data for the LLM and the human–LLM labels.
 Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$
 FAIDSet LLM-DetectAIve 0.7596 0.7697 0.7690 0.7653
-SVM (N-grams) 0.6764 0.6515 0.6126 0.6312
 T5-Sentinel 0.7568 0.7985 0.7840 0.7837
+SeqXGPT 0.6941 0.6802 0.6420 0.6603
 FAID 0.7964 0.8328 0.8352 0.8327
 LLM-DetectAIve LLM-DetectAIve 0.9049 0.9064 0.8352 0.8693
-SVM (N-grams) 0.8567 0.6516 0.6127 0.6313
 T5-Sentinel 0.8154 0.8137 0.8009 0.8105
+SeqXGPT 0.8712 0.7890 0.7608 0.7441
 FAID 0.9089 0.8817 0.8672 0.8737
 HART LLM-DetectAIve 0.8900 0.8787 0.8674 0.8715
-SVM (N-grams) 0.6347 0.5454 0.4385 0.4806
 T5-Sentinel 0.7852 0.7713 0.7834 0.7759
+SeqXGPT 0.6470 0.5582 0.4540 0.5075
 FAID 0.8996 0.9157 0.8648 0.8667
@@
 LLM-DetectAIve:
-Abassy et al. (2024) fine-tuned a RoBERTa Liu et al. (2019a) on fine-grained English AI-generated texts with learning rate of $2\times 10^{-5}$ for 10 epochs.
-Support Vector Machine with N-grams:
-Prova (2024) utilized SVM with N-gram features for binary AI text detection (Human vs.
-AI).
-We adapted this approach to a three-class detection.
-We first generated word-level trigram features using Count Vectorizer, and then used a Linear SVM to train with the gold labels to guide the classification.
+We adapted the method of Abassy et al.
+(2024) by fine-tuning a roberta-base sequence classification model.
+The input texts were tokenized using the RoBERTa tokenizer with a maximum sequence length of 256, and the model was trained for three-class detection.
 T5-Sentinel:
-Chen et al.
-(2023) utilized T5-Sentinel to classify text into two labels (Human vs.
-AI).
-This approach leverages the T5 model to predict the conditional probability of the next token, integrating the classification task into a sequence-to-sequence completion framework.
-This unique method directly leverages the strength of encoder-decoder model, which are trained to understand context and generate coherent text.
-5.3 Human, AI or Human-AI?
-Table 1 shows the performance of FAID and three baselines in four evaluation settings.
-FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domain, (iii) unseen generators, and (iv) unseen domain & generators settings, followed by LLM-DetectAIve in (i) and (iv), and T5-Sentinel in (ii) and (iii).
-Taking advantage of pretrained language models (i.e., RoBERTa, T5), they outperform SVM in most cases, particularly on HART dataset.
-FAID’s approach enhances the generalization performance over unseen domains and generators, compared with other methods.
-From the absolute accuracy, we can find that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55% respectively.
-The result suggests that FAID is an effective method to address the multilingual fine-grained AI-generated text detection task, improving the performance by leveraging multi-level contrastive learning to capture generalizable stylistic differences tied to AI families, rather than overfitting to surface-level artifacts.
+We adapt the T5-Sentinel framework introduced by Chen et al.
+(2023) for our three-class setup.
+Following the original configuration, we train using the AdamW optimizer (batch size 128, learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-5}$).
+This allows direct comparison with prior T5-based detectors under our experimental conditions.
+SeqXGPT:
+Following Wang et al.
+(2023) by modeling token-level likelihood patterns from LLM tokenizers for sentence-level detection, we adapted their method by updating the tokenizer models to align with our dataset’s label space.
+This adjustment ensures that the extracted token-level log-probability features correspond more accurately to the model types represented in our data.
+5.3 Human, LLM or Human-LLM?
+Table 1 shows the performance of FAID and three baselines in three evaluation settings.
+FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domains, (iii) unseen generators, and (iv) unseen domain & generators settings, followed by LLM-DetectAIve in (i) and (iv), and T5-Sentinel in (ii) and (iii).
+Although SeqXGPT aimed at extracting sequence-level features, it struggles with texts from advanced models, whose coherent and human-like writing styles reduce detectable distinctions.
+FAID improves generalization performance over unseen domains and generators compared to others.
+We can see that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55%, respectively.
+These results suggest that FAID is an effective method for addressing the multilingual fine-grained LLM-generated text detection task, improving performance by leveraging multi-level contrastive learning to capture generalizable stylistic differences tied to LLM families, rather than overfitting to surface-level artifacts.
 5.4 Identifying Different Generators
-We further examine the detector to identify specific AI generators which were seen during training.
-The goal of FAID is not only to detect AI involvement but also to identify the specific AI model or family applied, treating them as distinct authors.
+We further examine the detector to identify specific AI generators seen during training.
+The goal of FAID is not only to detect AI involvement, but also to identify the specific LLM or family applied, treating them as distinct authors.
 As shown in Table 2, across the three datasets, FAID consistently achieves higher performance compared to other baselines in almost all metrics.
@@
 FAID’s high performance when dealing with text originating from diverse known generators within these datasets indicates that FAID learned unique writing patterns and features of different AI generators by leveraging multi-level contrastive learning.
-6 Conclusion and Future Works
-In this work, we collected a multilingual fine-grained AI-generated text detection dataset FAIDSet with $\sim$84,000 examples.
-We further introduced FAID, a fine-grained AI-generated text detection framework aiming to distinguish between human-written, AI-generated, and human-AI collaborative texts.
-It combines multi-level contrastive learning and multi-task auxiliary objectives, with the perspective of treating AI model families as stylistic authors.
-FAID captures subtle writing patterns that generalize effectively across domains and evolving AI systems, where Fuzzy KNN-based inference method and training-free incremental adaptation enhance the robustness, enabling strong performance on unseen data.
-Experiments demonstrate that FAID consistently outperforms other strong baselines across multiple datasets and configurations.
-Its capability to recognize nuanced forms of collaborative texts and the ability to adapt to new generative models make it particularly suited for real-world applications, improving transparency in AI-assisted writing.
-In the future, we plan to expand FAIDSet by incorporating more languages, generators and domains, especially low-resource languages and informal text genres such as social media content or student essays, to further enhance cross-lingual and domain robustness.
-We also continue to investigate to deal with AI-generated text or more complex human-AI collaborative texts (i.e., from two or more AI families, generated by multiple AI families then edited by human) to deal with human’s behaviors in using AI for writing.
+6 Conclusion and Future Work
+We presented FAIDSet, a multilingual fine-grained LLM-generated text detection dataset comprising 84k examples, and FAID, a framework designed to distinguish between human-written, LLM-generated, and human–LLM collaborative texts.
+FAID integrates multi-level contrastive learning and multi-task auxiliary objectives, treating LLM families as stylistic “authors,” which enables it to capture subtle linguistic and stylistic cues that generalize effectively across domains and evolving generative systems.
+Moreover, its Fuzzy KNN-based inference and training-free incremental adaptation contribute to strong robustness and adaptability to unseen data.
+Our experiments demonstrated that FAID consistently outperforms competitive baselines across multiple datasets and settings.
+Its ability to detect nuanced collaborative writing and to adapt to emerging generative models highlights its potential for real-world deployment.
+In the future, we plan to extend FAIDSet to cover more languages, generators, and domains, particularly informal genres such as social media and student writing, to enhance cross-lingual and domain robustness further.
+We also aim to incorporate adversarial LLM-generated texts and more complex forms of human–LLM collaboration to capture the evolving dynamics of AI-assisted text creation.
 7 Limitations
-While FAID demonstrates strong performance and generalization across various domains and AI models, several limitations remain.
+While FAID demonstrates strong performance and generalization across various domains and LLMs, several limitations remain.
 First, although our dataset is multilingual and multi-domain, it is still limited in terms of low-resource languages and niche writing domains, which may affect performance in those contexts.
-Second, our framework is based on the fact that texts produced by AI models from the same family share similar stylistic features.
-However, this assumption may break down when a single text is influenced by multiple AI systems – such as when a human uses different models for drafting, rewriting, and polishing.
-In such cases, the resulting style may blend traits from multiple sources, making it more difficult to attribute authorship to a single AI family or clearly distinguish collaboration boundaries.
-8 Ethics and Broader Impacts
+FAIDSet is synthetic by construction.
+The controlled generation enables causal-style analysis, but it under-represents the messy, tool-chain-specific edits, wich are seen in the wild.
+We mitigate this with diverse prompt paraphrases, manual spot-checks, and generalization tests to held-out generators.
+Second, our framework is based on the observation that texts produced by LLMs from the same family share similar stylistic features.
+However, this may break down when a single text is influenced by multiple LLM systems – such as when a human uses different models for drafting, rewriting, and polishing a text.
+In such cases, the resulting style may blend traits from multiple LLMs, making it more difficult to attribute authorship to a single LLM family or clearly distinguish collaboration boundaries.
+8 Ethics and Broader Impact
 Data Collection and Licenses
 A primary ethical consideration is the data license.
-We reused existing dataset for our research: LLM-DetectAIve, HART and IELTS Writing, which have been publicly released under clear licensing agreements.
+We reused existing datasets for our research: LLM-DetectAIve, HART, and IELTS Writing, which have been publicly released under clear licensing agreements.
 We adhere to the intended usage of these dataset licenses.
 Security Implications.
-FAIDSet streamlines both the creation and rigorous testing of FAID.
-By spotting AI-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors.
+FAIDSet streamlines both the creation and the rigorous testing of FAID.
+By spotting LLM-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors.
 More broadly, it supports efforts to prevent the misuse of generative technologies in areas such as credential falsification.
-Detecting AI-crafted content across different languages can be tricky, due to language’s unique grammar and style.
+Detecting LLM-generated content across different languages can be tricky, due to language’s grammar and style.
 By enabling robust, multilingual and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly.
@@
 Bias and Fairness.
-AI-generated text detection systems may inadvertently encode or amplify biases present in training data.
+AI-generated text detection systems may inadvertently encode or amplify biases present in the training data.
 FAIDSet has been carefully constructed to include diverse domains and languages to reduce such biases.
@@
 Transparency and Reproducibility.
-In this work, to promote open research and community contributions, we published our code and data to the public.
+We promote open research and community contributions, and thus we publish our code and data.
 Appendix
 Appendix A Some Current Datasets on AI-generated Texts
-We do analysis in many current dataset on AI-generated text and summarized them in Table 3.
-Many dataset contains fine-grained labels and multi-generator, but all of analyzed datasets are monolingual – only in English.
-This problem makes us to collect and formulate a new dataset for multilingual, multi-domain and multilingual AI-generated text dataset to adapt the approach to real-world scenarios.
+We analyze many current datasets on AI-generated text and summarize them in Table 3.
+Many datasets contain fine-grained labels and multi-generators, but all of the analyzed datasets are monolingual – only in English.
+This problem prompts us to collect and formulate a new dataset for multilingual, multi-domain, and multilingual AI-generated text, enabling us to adapt the approach to real-world scenarios.
 Dataset Languages Label Space Domains Generators Size
@@
 Human-written, AI-polished
-Human-initiated, AI-continued
+Human-initiated, -continued
 AI-written, human-edited
@@
 GPT-3.5-turbo
-GPT-4o
+GPT‑4o
 Claude 3.5 Sonnet
@@
 67,000
-CyberHumanAI
-Najjar et al. (2025)
-English Human-written Wikipedia 500
 Table 3: English Fine-grained AI-generated Text Detection Datasets Overview.
@@
 B.1 Statistics
-Our FAIDSet includes 83,350 examples, which are separated into three subsets: train, validation, and test with the ratio shown in Table 4.
+Our FAIDSet includes 83,350 examples, which are divided into three subsets: train, validation, and test, with the ratios shown in Table 4.
 The data set also comprises various sources of human-written texts, which are described in Table 5.
-Subset Human AI Human-AI collab.
+Subset Human AI human–LLM collab.
 Train 14,176 12,076 32,091
@@
 Test 6,879 2,599 3,038
-Table 4: Number of examples per label in subsets in FAIDSet dataset.
+Table 4: Number of examples per label in subsets in the FAIDSet dataset.
 Source No. Human texts
@@
 B.2 Diverse Prompt Strategies
-To avoid biasing our generated corpora toward any single style or topic, we employ a broad set of prompt templates when synthesizing AI-generated and human-AI collaborative texts.
-By varying prompt structure, content domain, and complexity, we ensure that the resulting outputs cover a wide distribution of writing patterns, vocabulary, and rhetorical devices.
+To avoid biasing our generated corpora toward any single style or topic, we employ a broad set of prompt templates when synthesizing LLM-generated and human–LLM collaborative texts.
+By varying prompt structures, content domain, and complexity, we ensure that the resulting outputs cover a wide distribution of writing patterns, vocabulary, and rhetorical devices.
 This diversity helps our detector generalize more effectively to real-world inputs.
-Concretely, we sample two out of five prompts for AI-generated texts and several categories of human-AI collaborative texts in Table 6:
-  • AI-polished: Texts that were originally written by a human and then lightly refined by an AI system to improve grammar, clarity, or fluency without altering the core content or intent.
-  • AI-continued: Texts where a human wrote an initial portion (e.g., a sentence or paragraph), and an AI generated a continuation that attempts to follow the original style, tone, and intent.
-  • AI-paraphrased: Texts that were originally written by a human and then reworded by an AI system to express the same meaning using different phrasing, possibly altering sentence structure or word choice while preserving the original message.
-By mixing prompts across these categories, we generate a balanced corpus that mitigates over-fitting to any one prompt pattern and better reflects the diversity of real user queries.
-Category Student Thesis Paper Abstract
-AI-generated You are an IT student who are writing the graduation thesis.
-In a formal academic tone, write this section of a thesis on       , clearly outlining the structure, then write the passage concisely.
-The original text:       .
-In clear, structured prose, draft the section for a thesis titled       , cite some related works you mentioned in the passage and highlighting the contribution.
-The original text:       .
-You are a computer scientist who are very familiar with abstract writing for your works based on the title.
+Concretely, we use five prompts for LLM-generated texts in Table 6 and several categories of human–LLM collaborative texts in Table 7, Table 8 and Table 9 consecutively:
+  • LLM-polished: Texts that a human initially wrote and then lightly refined by an LLM system to improve grammar, clarity, or fluency without altering the core content or intent.
+  • LLM-continued: Texts where a human wrote an initial portion (e.g., a sentence or paragraph), and an LLM generated a continuation that attempts to follow the original style, tone, and intent.
+  • LLM-paraphrased: Texts that were initially written by a human and then reworded by an LLM system to express the same meaning using different phrasing, possibly altering sentence structure or word choice while preserving the original message.
+By mixing prompts across these categories, we generate a balanced corpus that mitigates over‐fitting to any one prompt pattern and better reflects the diversity of real user queries.
+Student Thesis Paper Abstract
+You are a university student majoring in computer science.
+Please briefly summarize the main idea of the following paragraph.
+After that, rewrite the paragraph based on this content.
+Write naturally in an academic style.
+The rewritten paragraph should be approximately the same length in characters as the original.
+The original text is:    In clear, structured prose, draft the section for a thesis titled   , cite some related works you mentioned in the passage, and highlight the contribution.
+The original text is:   .
+Please begin by briefly summarizing the main idea of the paragraph to ensure full comprehension and retention of all essential content.
+Then, rewrite the paragraph in a formal academic style, consistent with a university-level thesis.
+The rewritten section should read naturally, be coherent within the context of an academic paper, and have approximately the same character length as the original.
+Assume the role of a senior software engineer.
+Your task is to process a paragraph from a computer science thesis using a two-step method.
+Firstly, you must summarize by deconstruction, which involves analyzing the original text and providing a structured summary by identifying its primary purpose, key input parameters, and main outcomes.
+Secondly, rewrite the paragraph based on the structured summary.
+The new version must be technically precise, unambiguous, and logically structured, making it easy for another engineer to understand.
+The original text is:   .
+You are a research scientist preparing a paper for a top-tier computer science conference.
+The original text is:   .
+For the following paragraph from a thesis draft, you have to summarize the core contribution, which can be earned by beginning with providing a concise, one-sentence summary that captures the main scientific contribution or key finding of the paragraph.
+Then, rewrite the paragraph for publication by using the summary as a guide, ensuring it is written in a formal, objective, and precise tone suitable for a peer-reviewed publication.
+Ensure the rewritten text is information-dense yet easy for a fellow researcher to follow.
+You are a university student majoring in computer science.
+You need to extract the main idea by writing a summary of the paragraph’s main idea.
+Rewrite the paragraph based on the content you have summarized.
+The rewritten text should be in a formal academic style, read naturally, be coherent, and have approximately the same character length as the original.
+The original paragraph is:    Assume the role of a researcher with experience in writing abstracts for scientific papers.
+Write a short paragraph of approximately 150-200 words based on the topic conveyed by the provided file name.
+Start directly with the topic, presenting it clearly, objectively, and in a concise academic style.
+Use correct spelling and grammar, and write in a scholarly tone.
+Topic name:    You are a computer scientist who is very familiar with abstract writing for your works, based on the title.
 Craft a concise word_count-word abstract for a paper titled   , summarizing the problem statement, methodology, key findings, and contributions.
@@
 Compose a word_count-word abstract for the paper   , ensuring it includes motivation, approach, results, and implications for future research.
+Act like a senior researcher acting as a peer reviewer.
+Your task is to analyze and then rewrite the provided abstract to improve its structure and clarity.
+Deconstruct the abstract: First, examine the original text and break it down into four key components: What is the core problem being addressed?; What is the proposed solution or methodology?; What were the key results of the experiments?; What is the main contribution or impact of this work?
+After that, you must reconstruct the information from the components using only the data from your deconstructed points, and synthesize a new, cohesive abstract of approximately 150-200 words.
+Original abstract:   .
+You are a research scientist specializing in the sub-field suggested by the paper’s title.
+Your task is to generate a plausible abstract based only on the title.
+Based on the title, first generate a bulleted list of the likely components this paper would cover: the specific problem it probably addresses, the methodology it might propose, the kind of results one would expect, and its potential impact or contribution to the field.
+After that, weave these hypothesized points into a compelling and professional abstract of 150-200 words.
+Write it as if you were the author, confidently presenting your work.
+Paper title:    You are a technical writer for a prestigious AI research blog.
+Your goal is to rewrite a standard academic abstract to make it more impactful and highlight its core breakthrough for a broader technical audience.
+First, read the original abstract and identify the single most important takeaway or the core breakthrough of the paper and summarize this in one sentence.
+Second, rewrite the abstract to be approximately 150-200 words in length.
+Start with a strong opening sentence that directly states the problem or the breakthrough you identified.
+Then, briefly explain the methodology and results, always connecting them back to why they are important.
+The tone should be highly professional but more engaging than a typical arXiv abstract.
+Preserve all technical terms and citations accurately.
+Paper title:   .
+Original abstract:   .
+Table 6: List of diverse prompt templates used to generate FAIDSet – Label: LLM-generated
+Student Thesis Paper Abstract
+You are a university student majoring in computer science who has been assigned the task of polishing the paragraph below, which is excerpted from an undergraduate thesis.
+Improve the paragraph to make it clearer, more coherent, and more precise, while maintaining the original author’s academic tone and writing style.
+Do not rephrase any reference materials, figure labels, or citations preserve them exactly as they appear in the original paragraph.
 The original text:   .
-AI-polished You are an IT student who are refining your work to make it more complete.
+You are a meticulous academic editor with a specialization in computer science theses.
+Your task is to polish the following paragraph for conciseness and impact.
+Focus on eliminating redundant words and phrases, replacing weak verb constructions with stronger, more active verbs, and ensuring each sentence contributes directly and efficiently to the paragraph’s central point.
+The core technical meaning, all specific terminology, citations, and references to figures must be preserved precisely.
+The original text:   .
+Assume you are an IT student who is refining your work to make it more complete.
 Refine the following section excerpt for grammar, clarity, and academic style while preserving its original meaning and terminology.
+Provide only the polished version, without any introductory or explanatory text.
 The original text:   .
-You are an IT student who are refining your work to make it more complete.
+Act as a PhD candidate reviewing a section of an undergraduate thesis.
+Your primary goal is to enhance the logical flow and argumentative coherence of the paragraph below.
+Revise the text to ensure that sentences connect seamlessly with clear transitions.
+The paragraph should build a coherent argument from the opening sentence to the conclusion.
+Do not introduce new information or alter the original technical content, terminology, or references.
+Your focus is solely on restructuring the existing information to create a stronger, more persuasive, and logical narrative.
+The original text:   .
+You are a fourth-year IT student who is refining your work to make it more complete.
 Enhance the academic tone, coherence, and logical flow of this thesis section without altering technical content.
 The original text:   .
-You are a computer scientist who are very familiar with abstract writing and refining the written abstract.
+You are a researcher who has been assigned the task of polishing the paragraph below, which is excerpted from an undergraduate thesis.
+Improve the paragraph to make it clearer, more coherent, and more precise, while maintaining the original author’s academic tone and writing style.
+Do not rephrase any reference materials, figure labels, or citations—preserve them exactly as they appear in the original paragraph.
+The original text:  .
+You are a scientist who is very familiar with abstract writing and refining the written abstract.
 Improve the coherence, precision, and formal tone of this draft abstract without introducing new content.
+Provide only the polished version, without any introductory or explanatory text.
 The original text:  .
 Improve the clarity and conciseness of this abstract paragraph while maintaining all original findings and terminology.
+Provide only the polished version, without any introductory or explanatory text.
 The original text:   .
-AI-continued You are an IT student who are writing your graduation thesis.
-Continue to write the section from this thesis excerpt for approximately word_count words, maintaining formal academic structure and style.
+You are a program chair for a leading academic conference, skilled at identifying impactful research.
+Your task is to polish the following arXiv abstract to maximize its impact and make its core contribution immediately apparent.
+Focus on sharpening the opening sentence to act as a compelling hook.
+Rephrase and reorder sentences as needed to clearly convey the main findings and highlight the significance of the work.
+All original terminology, data, and citations must be strictly preserved.
+Provide only the polished version as a single, continuous paragraph.
 The original text:   .
-You are an IT student who are writing your graduation thesis.
+You are a meticulous editor for a top-tier scientific journal, such as Nature or Science.
+Your objective is to polish the following arXiv abstract to enhance its technical precision and information density.
+Scrutinize every word to ensure it is the most accurate choice.
+Refine phrasing to eliminate any ambiguity and, where supported by the text, replace qualitative descriptions with more specific, quantitative statements.
+The goal is a text where every clause delivers critical information efficiently.
+Do not alter the scientific findings, technical terms, or citations.
+The original text:   .
+Table 7: List of diverse prompt templates used to generate FAIDSet – Label: LLM-polished
+Student Thesis Paper Abstract
+You are a university student majoring in computer science who has been assigned the task of continuing the content of the paragraph below, which is excerpted from an undergraduate thesis.
+Please continue the text naturally, striving to mimic the tone and writing style of the given paragraph to avoid any inconsistency in expression, while ensuring clarity and coherence in an academic style.
+Do not rephrase the reference materials, figure labels, or citations-preserve them exactly as they appear in the original paragraph.
+The original text:   .
+You are an IT student who is writing your graduation thesis.
+Continue to write the section from file name    thesis excerpt for approximately word_count words, maintaining formal academic structure and style.
+Do not rephrase the reference materials, figure labels, or citations—preserve them exactly as they appear in the original paragraph.
+The original text:   .
+Act like an IT student who is writing your graduation thesis.
 Extend the section by adding supporting detailed information for a thesis on   .
+Do not rephrase the reference materials, figure labels, or citations—preserve them exactly as they appear in the original paragraph.
 The original text:   .
-You are a computer scientist who are very familiar with abstract writing.
-Add some concise concluding sentences to this partial abstract that highlights implications for future research.
+You are a computer science researcher meticulously documenting your work.
+Your task is to continue the paragraph starting with the sentence below.
+The continuation must elaborate on the underlying mechanism, process, or rationale implied by the initial statement, effectively answering the how or why.
+The completed paragraph should be logically sound and consistent with the topic of a thesis titled   .
+Maintain a formal academic tone and provide only the continuation as a single, seamless paragraph.
+Initial sentence:   .
+You are a final-year IT student analyzing your research findings for your graduation thesis.
+Continue the paragraph that begins with the key statement below by providing an analytical extension.
+Your writing should focus on comparing the statement to existing work, contrasting it with alternative approaches, or discussing its broader implications within the context of the thesis titled   .
+Ensure the analysis is coherent and maintains a scholarly tone.
+Initial statement:   .
+You are a researcher who has been assigned the task of continuing the content of the paragraph below, which is excerpted from an undergraduate thesis.
+Please continue the text naturally, striving to mimic the tone and writing style of the given paragraph to avoid any inconsistency in expression, while ensuring clarity and coherence in an academic style.
+Do not rephrase the reference materials, figure labels, or citations-preserve them exactly as they appear in the original paragraph.
+The original text:    You are a scientist who is very familiar with abstract writing.
+Add some concise concluding sentences to this partial abstract that highlight implications for future research.
+Do not rephrase the reference materials, figure labels, or citations-preserve them exactly as they appear in the original paragraph.
 The original text:   .
 Continue the abstract by writing a closing statement that underscores the study’s contributions and potential applications.
+Do not rephrase the reference materials, figure labels, or citations-preserve them exactly as they appear in the original paragraph.
 The original text:   .
-AI-paraphrased Paraphrase the following thesis section in a clear academic tone, preserving citations and technical terms exactly.
+You are a research scientist continuing the draft of a paper’s abstract.
+The provided text introduces the core problem or context.
+Your task is to continue the abstract by providing a concise description of the proposed methodology or approach.
+Detail the key techniques, model architecture, or experimental setup used to address the problem, ensuring the description is plausible for a paper titled   .
+The continuation must seamlessly connect to the initial text to form a single, coherent paragraph.
+Provide only the new text.
+Paper title:   , initial text:   .
+You are the lead author of a scientific paper summarizing your work.
+The text below already outlines the problem and methodology.
+Your task is to continue the abstract by presenting the key results and findings.
+Report on the primary outcomes, important performance metrics, or significant observations derived from your experiments.
+The results must be specific, quantitative where possible, and logically follow from the described method.
+The output must integrate smoothly with the initial text to form a single, cohesive paragraph.
+Initial text:   .
+Table 8: List of diverse prompt templates used to generate FAIDSet – Label: LLM-continued
+Student Thesis Paper Abstract
+You are an academic writing tutor.
+Your task is to perform a deep paraphrase of the following thesis excerpt.
+The goal is to create a version with significant structural and lexical differences from the original, while rigorously preserving the precise meaning, nuance, and all technical information.
+Focus on altering sentence construction and rephrasing ideas in a completely fresh way.
+All specific technical terms, citations, and figure labels must remain unchanged.
+The tone must remain formal and scholarly.
+The original text:   .
+You are a Computer Science Student tasked with paraphrasing the following paragraph, which has been extracted from the thesis of an undergraduate student.
+Paraphrase the given thesis content while preserving its original meaning and context.
+Maintain clarity, coherence, and an academic tone.
+Do not paraphrase: References, figure labels, and citations should remain unchanged.
+The original text:   .
+Paraphrase the following thesis section in a clear academic tone, preserving citations and technical terms exactly.
+Do not paraphrase: References, figure labels, and citations should remain unchanged.
+The original text:   .
+You are a senior researcher mentoring a student on their thesis.
+Paraphrase the following paragraph with the primary goal of improving clarity and directness.
+Untangle convoluted sentences and rephrase the content using a more straightforward structure.
+The aim is to express the same technical information in a way that is easier for a reader to parse, without losing any nuance or academic rigor.
+Do not alter or rephrase technical terminology, citations, or references to figures.
 The original text:   .
 Reword this thesis excerpt to improve readability and maintain its scholarly voice, keeping all references unchanged.
+Do not paraphrase: References, figure labels, and citations should remain unchanged.
 The original text:   .
-Rephrase this abstract in formal academic English, maintaining all original citations and technical accuracy.
+You are a researcher with paraphrasing the following paragraph, which has been extracted from the abstract of a science paper in the computer science domain.
+Paraphrase the given abstract content while preserving its original meaning and context.
+Maintain clarity, coherence, and an academic tone.
+Do not paraphrase: References, figure labels, and citations should remain unchanged.
+The original text:    Rephrase this abstract in formal academic English, maintaining all original citations and technical accuracy.
+Do not paraphrase: References, figure labels, and citations should remain unchanged.
+The original text:   .
+You are an expert scientific editor tasked with reframing an abstract to maximize its immediate impact.
+Perform a structural paraphrase on the following text.
+First, identify the core components of the abstract (Problem, Method, Results, Contribution) internally.
+Then, rephrase and reorder these components to lead with the main Contribution or Result, followed by the problem it solves and the method used.
+This inverted structure should create a fresh and compelling narrative while preserving all original information.
+Strict requirements: All technical terms, data, and citations must be preserved exactly.
 The original text:   .
 Paraphrase this abstract paragraph to enhance clarity and flow, ensuring all technical terms and citations remain intact.
+Do not paraphrase: References, figure labels, and citations should remain unchanged.
 The original text:   .
-Table 6: Samples of diverse prompt templates used to generate FAIDSet
-Appendix C Analysis on similarity across models and model families
-To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis using various perspectives: N-grams distribution, text length patterns, and semantic embedding visualization.
-These analyses allow us to examine similarities both within the same model family and across different model families, leading to a robust understanding of LLM "authorship" characteristics.
-C.1 Text Distribution Between AI Families
-We first examined the distribution of text length, measured in both word and character counts across outputs from five AI models: Llama-3.3-70B-Instruct-Turbo Dubey et al.
+You are a senior scientist adapting a specialized abstract for a broader scientific audience.
+Your task is to paraphrase the following text to make it more accessible to researchers in adjacent fields, without sacrificing technical rigor.
+Rephrase the abstract by substituting hyper-specific jargon with more widely understood technical equivalents, but only if the precise meaning is retained.
+The goal is for a researcher from a different subfield to quickly grasp the core concepts.
+Strict requirements: The abstract’s original meaning, key findings, and data must be perfectly preserved.
+All citations and figure references must remain unchanged.
+The original text:
+Table 9: List of diverse prompt templates used to generate FAIDSet – Label: LLM-paraphrased
+Appendix C Experimental Details and Real-World Use Cases
+C.1 Experimental Setup
+Unless noted otherwise, we use a batch size of 64, the AdamW optimizer with a learning rate of 2e52e-5 and weight decay of 1e41e-4, 50 epochs, and 2000 warm-up steps.
+For the fuzzy kNN component, we set top-K = 20.
+We use temperature = 0.7.
+C.2 Computational Cost
+All experiments run on a single NVIDIA A100 (40 GB); total wall-clock training time for FAID on FAIDSet is approximately 5 hours.
+In our setup using FAISS Douze et al.
+(2024) with CPU inference, the average query latency is approximately 5ms for FAIDSet with 84,000 embeddings on a standard server (Intel Xeon CPU, 64GB RAM).
+When performing with Fuzzy KNN, it takes some minor time to process, so the total average time for inference is 10ms.
+C.3 Real-world Management of Vector Database
+In the case of FAIDSet, our entire training and validation embedding store requires approximately 200MB, which is easily handled by standard server disks.
+For larger datasets, we believe the system remains scalable:
+  • FAISS supports disk-based indices and optimized search methods to keep memory usage low even with millions of vectors.
+  • Modern servers with hundreds of GB of disk are sufficient for storing large embedding banks, and the use of CPU-only inference (since Fuzzy KNN is GPU-free) makes the architecture cost-effective for deployment in resource-constrained environments.
+C.4 Generalizability of FAID on unseen human–LLM type of collaboration
+To further assess FAID’s robustness to unseen collaborative writing patterns between humans and LLM systems, we conducted an additional experiment focusing on hybrid authorship styles that were not present in the original training dataset.
+We performed a manual quality control process to enhance the reliability of our dataset and ensure the robustness of the FAID model.
+A team of five annotators, all IT-major university students fluent in English and aged 18-25, participated in the annotation process.
+Each annotator was assigned approximately 80 samples, covering all collaborative writing styles (LLM-generated, LLM-continued,LLM-paraphrased, and LLM-polished), resulting in a total of 400 human-reviewed instances representing various collaboration modes.
+During the manual revision stage, annotators followed two key quality control principles:
+  1. 1.
+    Ensuring logical and informational consistency between the output samples and original texts, where the output length was controlled to be between 70%–150% of the source length.
+  2. 2.
+    Improving linguistic quality through spelling correction, synonym replacement, and contextual word refinement to ensure natural fluency and stylistic coherence.
+FAID was evaluated on this manually curated dataset without any further fine-tuning.
+Despite these samples representing an unseen distribution, the model achieved a strong overall accuracy of 84.8%, precision of 82.8%, and recall of 85.0%, across all collaboration categories.
+This suggests a strong generalization capability beyond the data distributions encountered during model development.
+It also highlights FAID’s sensitivity to fine-grained stylistic blending, where human revision only partially conceals the generative footprint of LLMs.
+C.5 Generalizability of FAID on real-world scenarios.
+To further assess the generalizability of FAID beyond the controlled benchmark setting, we conducted an additional user study simulating realistic academic and professional writing scenarios.
+The goal was to evaluate whether FAID can maintain detection accuracy when applied to authentic, unconstrained text produced through real interactions with LLMs.
+A group of five volunteers with diverse academic backgrounds (same as in the Appendix C.4).
+Each participant was instructed to engage with four popular AI systems: ChatGPT, Gemini, DeepSeek, and Llama 3.1 on Together.AI to generate text resembling authentic student or professional writing tasks.
+The interactions were open-ended but guided by three types of prompts commonly encountered in real-world use cases: writing a paper summary (as an abstract) and writing a passage for their own graduation thesis.
+Each participant was asked to generate five outputs per model, yielding a total of 200 real-world text samples (5 participants $\times$ 4 models $\times$ 10 trials).
+To ensure authenticity and diversity, participants were encouraged to adjust prompts iteratively, include follow-up clarifications, and perform light editing, mimicking realistic human–LLM co-writing behavior.
+FAID achieved strong performance in this real-world setting, with an overall accuracy of 88.5%, precision of 85.9%, and recall of 89.7%, indicating that FAID still exhibits strong generalizability in uncontrolled real-world scenarios with in-domain settings.
+Appendix D Analysis on Similarity across Models and Model Families
+To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis from various perspectives, including N-gram distribution, text length patterns, and semantic embedding visualization.
+These analyses allow us to examine similarities both within the same model family and across different model families, leading to a robust understanding of AI "authorship" characteristics.
+D.1 Text Distribution between LLM Families
+We first examined the distribution of text length, measured in both word and character counts across outputs from five LLMs: Llama-3.3-70B-Instruct-Turbo Dubey et al.
 (2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Team et al.
@@
 This reinforces the hypothesis that text length behavior is not only model-dependent but also family-coherent, with Gemini models forming a distinct cluster.
-C.2 Text Distribution Between AI Models Within the Same Family
-We performed N-grams frequency analysis on responses generated by three models within the Gemini family: Gemini 2.0, Gemini 2.0 Flash, and Gemini 1.5 Flash using 500 texts from the arXiv abstract dataset.
+D.2 Text Distribution Between LLMs Within the Same Family
+We performed N-gram frequency analysis on responses generated by three models within the Gemini family: Gemini 2.0, Gemini 2.0 Flash, and Gemini 1.5 Flash using 500 texts from the arXiv abstract dataset.
 Figure 5 highlights overlapping high-frequency tokens and similar patterns in word usage and phrase structure among the three models.
 Despite minor differences in architectural speed (e.g., Flash vs.
-regular) or release chronology, the N-grams distributions show minimal divergence.
+regular) or release chronology, the N-gram distributions show minimal divergence.
 Frequently used tokens, such as domain-specific terms and transitional phrases, appeared with nearly identical frequencies.
@@
 These intra-family similarities support treating model variants within a family as a unified authoring entity when performing analysis or authorship attribution.
-C.3 Embedding Visualization and Semantic Cohesion
-To delve into deeper semantic alignment, we visualized embeddings generated by unsupervised SimCSE XLM-RoBERTa-base model of generated texts from two model families: Gemini family (texts generated by Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash), and GPT-4o/4o-mini using PCA to reduce the high-dimensional embeddings into a space.
+D.3 Embedding Visualization and Semantic Cohesion
+To delve into deeper semantic alignment, we visualized embeddings generated by an unsupervised SimCSE XLM-RoBERTa-base model of generated texts from two model families: Gemini family (texts generated by Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash), and GPT-4o/4o-mini using PCA to reduce the high-dimensional embeddings into a space.
 As shown in Figure 4, Gemini model embeddings form tight, overlapping clusters, indicating a high degree of semantic cohesion among their outputs.
@@
 This visualization confirms that the Gemini family not only shares stylistic features but also maintains semantic coherence, distinguishing it from models of other families at a conceptual level.
-C.4 Conclusion
-The consistency observed across N-grams distributions, text length patterns, and semantic embeddings among Gemini models substantiates our decision to treat each AI family as an author.
-These models demonstrate coherent writing styles, shared lexical preferences, and tightly clustered semantic representations - hallmarks of unified authorship.
-Conversely, inter-family comparisons show clear separability, emphasizing the distinctiveness of each AI model family’s writing behavior.
-(a) Word length distribution across five AI models.
-(b) Character length distribution across five AI models.
+D.4 Conclusion
+The consistency observed across N-gram distributions, text length patterns, and semantic embeddings among Gemini models substantiates our decision to treat each LLM family as an author.
+These models demonstrate coherent writing styles, shared lexical preferences, and tightly clustered semantic representations, hallmarks of unified authorship.
+Conversely, inter-family comparisons show clear separability, emphasizing the distinctiveness of each LLM family’s writing behavior.
+(a) Word length distribution across five LLMs.
+(b) Character length distribution across five LLMs.
 Figure 3: Text length distributions in words and characters across Llama-3.3, GPT-4o/4o-mini, Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash.
@@
 (c) Top 3-grams of Gemini 1.5 Flash (500 samples).
-Figure 5: Top 20 most common 3-gram from Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash using 500 sample prompts.
-Appendix D Model selection for detector
+Figure 5: Top 20 most common trigrams from Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash using 500 sample prompts.
+Appendix E Model Selection for Detector
 Model #. Params Known Generators Unseen Generators
@@
 Base model comparison.
-We first evaluated three popular monolingual models: RoBERTa-base, Flan-T5-base, and e5-base-v2 using the same training and evaluation splits.
-RoBERTa-base Liu et al.
-(2019b) and e5-base-v2 Wang et al.
-(2022) achieved the most balanced trade-off between classification accuracy and regression error (MSE, MAE), while Flan-T5-base Chung et al.
-(2022) lagged slightly in F1-macro.
-These results indicated that a stronger encoder backbone yields more robust performance and motivated exploration of multilingual variants for further gains.
+We first evaluated three popular monolingual models: RoBERTa‑base, Flan‑T5‑base, and e5‑base‑v2 using the same training and evaluation splits.
+RoBERTa‑base Liu et al.
+(2019) and e5‑base‑v2 Wang et al.
+(2022) achieved the most balanced trade‑off between classification accuracy and regression error (MSE, MAE), while Flan‑T5‑base Chung et al.
+(2022) lagged slightly in F1‑macro.
+These results indicate that a stronger encoder backbone yields more robust performance, motivating the exploration of multilingual variants for further gains.
 Multilingual variants.
-Next, we experimented with XLM-RoBERTa-base Conneau et al.
-(2019) and Multilingual-e5-base Wang et al.
+Next, we experimented with XLM‑RoBERTa‑base Conneau et al.
+(2019) and Multilingual‑e5‑base Wang et al.
 (2024a).
-Both models leverage cross-lingual pretraining, which in our scenario appears to improve representation of diverse linguistic patterns present in FAIDSet.
-In particular, XLM-RoBERTa-base delivered a substantial boost in all metrics, suggesting that its multilingual training enhances generalization even on primarily monolingual inputs.
+Both models leverage cross-lingual pretraining, which, in our scenario, appears to improve the representation of diverse linguistic patterns present in FAIDSet.
+In particular, XLM‑RoBERTa‑base delivered a substantial boost in all metrics, suggesting that its multilingual training enhances generalization even on primarily monolingual inputs.
 Contrastive learning with SimCSE.
@@
 (2021) to further refine the sentence embeddings.
-We tested both supervised (trained on NLI data) and unsupervised (trained on Wikipedia corpus) SimCSE variants applied to RoBERTa-base.
+We tested both supervised (trained on NLI data) and unsupervised (trained on Wikipedia corpus) SimCSE variants applied to RoBERTa‑base.
 The unsupervised version outperformed its supervised counterpart, confirming prior findings that unsupervised SimCSE often yields stronger semantic encoders for downstream tasks.
-Encouraged by this, we adopted the unsupervised SimCSE approach on XLM-RoBERTa-base Vamvas and Sennrich (2023), which achieved the highest accuracy and lowest error rates across all configurations.
-Based on these experiments, we selected the unsupervised SimCSE XLM-RoBERTa-base model for our final system.
+Encouraged by this, we adopted the unsupervised SimCSE approach on XLM‑RoBERTa‑base Vamvas and Sennrich (2023), which achieved the highest accuracy and lowest error rates across all configurations.
+Based on these experiments, we selected the unsupervised SimCSE XLM‑RoBERTa‑base model for our final system.
 Algorithm Known Generators Unseen Generators
@@
 Table 11: Comparison of clustering algorithms on known vs. unseen generators. The best results are shown in bold.
-Appendix E Ablation study
-E.1 The need to use a vector database
+Appendix F Ablation Study
+F.1 The Need to Use a Vector Database
 We first used the trained model to classify the text and observed a significant performance decline from 90% when evaluating on unseen domains or text generated by novel models.
@@
   • Generator-Independent Coverage: As novel text generators emerge, their embeddings populate the database; retrieval naturally adapts to new styles or patterns by finding the closest existing vectors.
-E.2 Clustering Algorithm Selection
+F.2 Clustering Algorithm Selection
 To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies on the encoded vectors stored in our vector database.
-Each algorithm was tasked with grouping text samples into human-written, AI-generated, and human-AI collaborative categories, using both known-generator data (held-out from training) and entirely unseen generator data.
+Each algorithm was tasked with grouping text samples into human‐written, AI‐generated, and human–LLM collaborative categories, using both known‐generator data (held out from training) and entirely unseen generator data.
 Performance was measured in terms of accuracy and F1-macro score.
-We encoded each text sample using our pretrained sequence-classification model’s penultimate layer, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three text-origin classes.
+We encoded each text sample using the penultimate layer of our pre-trained sequence-classification model, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three text-origin classes.
 From the results in Table 11, we draw some key findings:
-  • Traditional algorithms show reasonable performance on held-out known generators but degrade notably on unseen generators.
-  • Fuzzy C-Means leverages membership degrees to handle overlapping distributions, improving both accuracy and F1-macro by about 4% over k-Means, with smaller degradation on unseen data.
+  • Traditional algorithms show reasonable performance on held‐out known generators but degrade notably on unseen generators.
+  • Fuzzy C‐Means leverages membership degrees to handle overlapping distributions, improving both accuracy and F1-macro by about 4% over k‐Means, with smaller degradation on unseen data.
   • Fuzzy KNN combines local neighbor information with fuzzy membership, achieving the best overall performance: 95.2% accuracy and 95.1% F1-macro on known generators, and 93.3% accuracy and 93.3% F1-macro on unseen generators.
-Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our full architecture.
+Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our whole architecture.

### v3

@@
 Abstract
-The growing collaboration between humans and AI models in generative tasks has introduced new challenges in distinguishing between human-written, LLM-generated, and human–LLM collaborative texts.
+The growing collaboration between humans and LLMs in generative tasks has introduced new challenges in distinguishing between human-written, LLM-generated, and human–LLM collaborative texts.
 In this work, we collect a multilingual, multi-domain, multi-generator dataset FAIDSet.
-We further introduce a fine-grained detection framework FAID to classify text into these three categories, and also to identify the underlying LLM family of the generator.
+We further introduce a fine-grained detection framework, FAID, to classify text into these three categories and to identify the underlying LLM family of the generator.
 Unlike existing binary classifiers, FAID is built to capture both authorship and model-specific characteristics.
 Our method combines multi-level contrastive learning with multi-task auxiliary classification to learn subtle stylistic cues.
-By modeling LLM families as distinct stylistic entities, we incorporate an adaptation to address distributional shifts without retraining for unseen data.
-Our experimental results demonstrate that FAID outperforms several baselines, particularly enhancing the generalization accuracy on unseen domains and new LLMs, thus offering a potential solution for improving transparency and accountability in AI-assisted writing.
+By modeling LLM families as distinct stylistic entities, we adapt to address distributional shifts without retraining on unseen data.
+Our results demonstrate that FAID outperforms several baselines, particularly improving generalization accuracy across unseen domains and new LLMs, offering a potential solution to improve transparency and accountability in AI-assisted writing.
+Our data and code are available at https://github.com/mbzuai-nlp/FAID
+**footnotetext: Equal contribution.
 1 Introduction
-LLMs have evolved from an assistant tool to a creator or initiator, from helping polish papers and generate code to initiate proposals and draft essays, while humans increasingly serve as optimizers and reviewers.
+LLMs have evolved from an assistant tool to a creator or initiator, from helping polish papers to initiating proposals and drafting essays, while humans increasingly serve as optimizers and reviewers.
 In such deeply collaborative human–LLM settings, measuring human contribution becomes challenging, yet clarifying authorship is critical for accountability and transparency, particularly in educational and academic contexts Wang et al.
@@
 LLM Wang et al.
-(2024d, b), or on fine-grained detection limited to English Abassy et al.
+(2024e, d); Su et al.
+(2023) or on fine-grained detection limited to English Abassy et al.
 (2024); Koike et al.
@@
 (2023); Zhang et al.
-(2024a).
+(2024).
 Moreover, both struggle with generalization to unseen domains, languages, and LLM generators Wang et al.
@@
 (2024).
-Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain and multi-generator dataset, FAIDSet, for fine-grained detection, i.e., identifying a text into three categories: LLM-generated, human-written, and human–LLM collaborative; and (ii) introducing a detection framework FAID to improve generalization performance.
-Our dataset FAIDSet focuses on the academic field including paper abstracts, student theses, and scholarly reports, and contains generation by a variety of families of LLMs, e.g., GPT, Gemini, DeepSeek, and Llama OpenAI (2024); Team et al.
-(2024); DeepSeek-AI (2024); Dubey et al.
+Our work aims to bridge this gap by (i) collecting a multilingual, multi-domain, and multi-generator dataset, FAIDSet, for fine-grained detection, i.e., identifying a text into three categories: LLM-generated, human-written, and human–LLM collaborative, and (ii) introducing a framework FAID to improve generalization performance.
+Our dataset FAIDSet [^1] https://huggingface.co/datasets/ngocminhta/FAIDSet focuses on the academic field, including paper abstracts, student theses, and reports, and contains generation by a variety of families of LLMs, e.g., GPT, Gemini, DeepSeek, and Llama OpenAI (2024); Gemini Team (2024); DeepSeek-AI (2025); Dubey et al.
 (2024).
-Our detection framework FAID treats each LLM as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features.
-We achieve this by optimizing a language encoder with multi-level author relationship (e.g., the stylistic similarity between texts from the same LLM family is greater than the that between a human and an LLM) to capture author-specific distinguishable signals of an input text using contrastive learning, along with the task of fine-tuning a classifier to recognize the input text’s origin.
-This multitask learning process forces the encoder to reorganize the hidden space, so that the representations of texts by the same authors are distributed closer than such by different authors.
-In practice, given that generations produced by LLMs from same company tend to have similar writing styles due to similar model architecture, training data, and training strategy (see Appendix D for more detail), we consider each LLM family as an “author.” This can also help the detector acquire prior knowledge of future LLMs from the same company.
+Our detection framework, FAID, treats each LLM as a distinct author, learning specific features in the hidden space to differentiate different authors, instead of classifying based on hand-crafted stylistic features.
+We achieve this by optimizing a language encoder with multi-level author relationship (e.g., the stylistic similarity between texts from the same LLM family is greater than that between a human and an LLM) to capture author-specific distinguishable signals of an input text using contrastive learning, along with the task of fine-tuning a classifier to recognize the input text’s origin.
+This multitask learning process forces the encoder to reorganize the hidden space so that the representations of texts by the same authors are distributed closer together than those by different authors.
+In practice, given that generations produced by LLMs from the same company tend to have similar writing styles due to similar model architecture, training data, and training strategy (see Appendix D for more detail), we consider each LLM family as an “author.” This can also help the detector acquire prior knowledge of future LLMs from the same company.
 Our experiments show that FAID consistently outperforms other baseline detectors in both in-domain and out-of-domain evaluations.
 Our contributions can be summarized as follows:
-  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained LLM-generated text detection with 84k examples.
-  • We propose a detection framework FAID to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in the hidden representation.
-  • We show that FAID outperforms baseline detectors, particularly over unseen domains and generators.
-Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to other texts from our database, labeled with ground truth authorship labels.
+  • We collect a new multilingual, multi-domain, multi-generator dataset for fine-grained LLM-generated text detection with 83,350 examples.
+  • We propose a detection framework, FAID, to improve generalization performance in unseen domains and generators by capturing subtle stylistic features in the hidden representation.
+  • We show that FAID outperforms baseline detectors, particularly in unseen domains and with unseen generators.
+Meanwhile, the nature of FAID allows us to assess the stylistic proximity of a given text to other texts in our database, each labeled with ground-truth authorship labels.
 2 Background and Related Work
 Advancements in LLMs have fundamentally reshaped the process of text production and refinement.
-Rather than originating every word independently, people increasingly assume the roles of post-editors or reviewers, intervening after an initial draft has been generated by an LLM.
+Rather than originating every word independently, people increasingly assume the role of post-editors or reviewers, intervening after an LLM has generated an initial draft.
 This collaborative paradigm extends across diverse domains, including academic publishing, journalism, education, and social media, thus making hybrid human–AI authorship an emerging norm Cheng et al.
@@
 (2022).
-Such a shift calls question into binary authorship detection: human vs.
+Such a shift calls into binary authorship detection: human vs.
 AI.
@@
 (2025).
-Addressing these concerns necessitate fine-granularity authorship detection, and even tracing back to a specific LLM family, in order to assess the extent of human contribution Hutson (2025).
-Ensuring transparent disclosure of LLM involvement is critical for upholding research honesty and preserving the integrity of research.
+Addressing these concerns requires fine-grained authorship detection, even tracing it back to a specific LLM family, to assess the extent of human contribution Hutson (2025).
+Ensuring transparent disclosure of LLM involvement is critical to upholding research integrity and honesty.
 In response to these challenges, we collect a new dataset, FAIDSet, and propose a detection framework, FAID, which generalizes well to new domains, languages, and models, achieving consistently high accuracy and reliability.
 2.1 Fine-Grained AI-generated Text Datasets
-Many prior studies have explored fine-grained AI-generated text datasets across various forms of human–AI collaboration, e.g., MixSet (Zhang et al., 2024b), and Beemo (Artemova et al., 2025).
-See Table 3 in the Appendix, for a discussion of more datasets and detailed information regarding the label space, and the coverage of domains, languages and LLMs for each one.
+Many prior studies have explored fine-grained AI-generated text datasets across various forms of human–AI collaboration, e.g., MixSet (Zhang et al., 2024) and Beemo (Artemova et al., 2025).
+See Appendix A for a discussion of more datasets and detailed information regarding the label space, and the coverage of domains, languages, and LLMs for each one.
 However, all these studies were limited to English.
 A substantial gap remains in the availability of large-scale, fine-grained multilingual LLM-generated text detection datasets.
-To bridge this gap, we collect a multilingual fine-grained LLM-generated text detection dataset, which encompasses 84k texts generated by the latest LLMs and includes diverse forms of human–LLM collaborative generations.
+To bridge this gap, we collect a multilingual fine-grained LLM-generated text detection dataset, which encompasses 83k texts generated by the latest LLMs and includes diverse forms of human–LLM collaborative generations.
 This dataset can facilitate the development of more robust and generalizable detection models that are capable of handling complex multilingual collaborative scenarios.
-2.2 Generalization of AI-Generated Text Detection
+2.2 Generalization of LLM-Generated Text Detection
 A recent study, M4GT-Bench, Wang et al.
 (2024c) has highlighted a persistent challenge for both binary and fine-grained AI-generated text detection: poor generalization to unseen domains, languages, and generators.
-Many detection methods have shown a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors to deal with real-world scenarios and evolving LLMs outputs.
+Many detection methods have shown a significant drop in performance on out-of-distribution data, underscoring the difficulty of building robust detectors for real-world scenarios and evolving LLM outputs.
 Various approaches have been proposed to improve generalization.
 OUTFOX Koike et al.
-(2024) leverages adversarial in-context learning, dynamically generating challenging examples to enhance robustness, but still faces limitations in domain transferability and computational efficiency.
+(2024) leveraged adversarial in-context learning to dynamically generate challenging examples that enhance robustness, but still faced limitations in domain transferability and computational efficiency.
 LLM-DetectAIve Abassy et al.
-(2024) adopts fine-grained classification and incorporates domain-adversarial training to reduce overfitting; yet, its generalization to unseen domains and generators remains limited; moreover, its current version lacks multilingual support.
+(2024) adopted fine-grained classification and incorporated domain-adversarial training to reduce overfitting; however, its generalization to unseen domains and generators remains limited, and its current version lacks multilingual support.
 SeqXGPT Wang et al.
-(2023) focuses on sentence-level detection by combining log-probabilities with convolutional and self-attention mechanisms, which helps to capture subtle mixed-content signals and improves generalization across input styles.
-However, its reliance on specific model features and limited semantic representation constrains adaptability to new generators and domains.
+(2023) focused on sentence-level detection by combining log-probabilities with convolutional and self-attention mechanisms, thereby helping capture subtle mixed-content signals and improving generalization across input styles.
+However, its reliance on specific model features and its limited semantic representation constrained the adaptability to new generators and domains.
 Finally, the DeTeCtive framework Guo et al.
-(2024), introduces multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, but it primarily focuses on binary classification and does not fully address human–AI collaborative texts.
+(2024) introduced multi-level contrastive learning to better capture writing-style diversity and enhance generalizability, especially for out-of-distribution scenarios, but it primarily focused on binary classification and did not fully address human–AI collaborative texts.
 2.3 Contrastive Learning for AI-Generated Text Detection
-Contrastive learning has been extensively used to improve sentence representation by pulling semantically similar sentences closer and pushing dissimilar ones apart.
+Contrastive learning has been widely used to improve sentence representations by pulling semantically similar sentences closer together and pushing dissimilar ones apart.
 SimCSE treats a sentence subjected to dropout noise as a semantically similar counterpart (i.e., a positive pair) and trains the encoder to minimize the distance between the original and the noise sentence.
@@
 DeCLUTR Giorgi et al.
-(2021) constructed positive pairs by extracting different spans from the same texts, and by sampling negative pairs from different texts.
-We adopt the same core idea, but reorganize the latent space based on writing style, pulling human-written texts to cluster together while remaining distant from LLM-generated ones.
-Similarly to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement in collaborative texts, ranging from 0 (fully LLM-generated) to 1 (fully human-written).
+(2021) constructed positive pairs by extracting different spans from the same texts and by sampling negative pairs from different texts.
+We adopt the same core idea, but reorganize the latent space by clustering human-written texts by writing style, keeping them distant from LLM-generated texts.
+Similarly to semantic textual similarity tasks, where sentence similarity ranges from 0 to 5 to reflect varying degrees of semantic overlap, we incorporate ordinal regression into our framework to model the degree of human involvement, ranging from 0 (solely LLM) to 1 (fully human).
 Another work, DeTeCtive Guo et al.
-(2024), also took advantage of contrastive learning and was used in binary AI-generated text detection.
-Based on a multi-task framework, it was trained to learn writing-style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs.
+(2024), also leveraged contrastive learning and was used for binary detection task.
+Based on a multi-task framework, it was trained to learn style diversity using a multi-level contrastive loss, and an auxiliary task of classifying the source of a given text (human vs.
 AI) to capture distinguishable signals.
-For inference, their pipeline encoded the input text as a hidden vector and applied dense retrieval to match the cluster, depending on the stylistic similarity with respect to a database of previously indexed training features.
-Additionally, instead of retraining the model when faced with new data, they encoded new data with the trained encoder to obtain embeddings, and then added them to the feature database to augment.
+For inference, their pipeline encoded the input text as a hidden vector and used dense retrieval to match the cluster based on stylistic similarity against a database of previously indexed training features.
+Additionally, instead of retraining the model on new data, they encoded the new data with the trained encoder to obtain embeddings, then added them to the feature database to augment.
 This largely improved the generalizability to unseen domains and new generators.
@@
 LLM-generated) while overlooking the increasingly prevalent class of human–LLM collaborative texts.
-Our work bridges this gap with the aim of enhancing generalization performance for fine-grained LLM-generated text detection.
+Our work bridges this gap to enhance generalization performance in fine-grained LLM-generated text detection.
 3 FAIDSet
-We collected a new multilingual, multi-domain and multi-generator LLM-generated text dataset, FAIDSet.
-It contains texts generated by LLMs, written by humans and collaborated by both humans and LLMs, resulting in a total of 84k examples; see Appendix B for more details about FAIDSet.
+We collected a new multilingual, multi-domain, and multi-generator LLM-generated text dataset, FAIDSet.
+It contains texts generated by LLMs, written by humans, and collaborated by both, resulting in a total of 83,350 examples; see Appendix B for more details about FAIDSet.
 FAIDSet covers two domains: student theses and paper abstracts, where identifying authorship is critical, across two languages (Vietnamese and English).
-We collected student theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^1] https://vjol.info.vn/ (VJOL).
+We collected students’ theses from the database of Hanoi University of Science and Technology, and paper abstracts from arXiv and Vietnam Journals Online [^2] https://vjol.info.vn/ (VJOL).
 Models and Label Space.
@@
 Diverse Prompt Strategies.
-We generated data with diverse tones and contexts while ensuring the accuracy of the content.
+We generated data with diverse tones and contexts while ensuring content accuracy.
 Depending on the data source and context, we crafted prompts to create varied outputs suitable for different real-world scenarios.
@@
 See Appendix B.2 for the full list.
-While FAIDSet does not capture the full natural diversity of in-the-wild LLMs outputs, the controlled setup enables us to systematically model stylistic and collaborative signals across multiple languages and families.
+While FAIDSet does not capture the full natural diversity of in-the-wild LLMs’ outputs, the controlled setup enables us to systematically model stylistic and collaborative signals across multiple languages and families.
 This makes FAIDSet both a reproducible training resource and a benchmarking corpus: it provides reliable supervision for developing detectors while also serving as a testbed for assessing generalization to more diverse, unseen scenarios.
 Quality Control.
-To avoid bad machine-generated texts, which can introduce remarkably distinguishable signals, we performed quality control by randomly sampling 10-20 instances for each domain, source, and LLM generator.
-Manual inspection focused on fluency, coherence and factual plausibility.
-Overall, the generated texts demonstrated high linguistic quality with most outputs being fluent and logically reasonable.
-Nevertheless, occasional issues such as repetitive phrasing, incomplete reasoning or overuse of formal expressions were observed.
-In that cases, we adjusted the prompts (e.g., by specifying desired length or style) or refined generation parameters to improve diversity and logical consistency.
+To avoid bad machine-generated texts, which can introduce remarkably distinguishable signals, we performed quality control by randomly sampling 10–20 instances for each domain, source, and LLM generator.
+Manual inspection focused on fluency, coherence, and factual plausibility.
+Overall, the generated texts demonstrated high linguistic quality, with most outputs being fluent and logically reasonable.
+Nevertheless, occasional issues such as repetitive phrasing, incomplete reasoning, or overuse of formal expressions were observed.
+In those cases, we adjusted the prompts (e.g., by specifying desired length or style) or refined generation parameters to improve diversity and logical consistency.
 After these adjustments, the quality across generators and domains was found to be stable and satisfactory.
@@
 Our task is a three-class classification problem: human-written vs. LLM-generated vs. human–LLM collaborative text detection.
-The human–LLM collaborative category involves a range of interaction between humans and LLM systems, such as human-written, LLM-polished; human-initiated, LLM-continued; human-written, LLM-paraphrased, etc.
+The human–LLM collaborative category involves a range of interactions between humans and LLM systems, such as (a) human-written, LLM-polished, (b) human-initiated, LLM-continued, (c) human-written, LLM-paraphrased, etc.
 Given the growing variety and complexity (e.g., deeply mixed text) of collaborative patterns, this is not exhaustive.
@@
 Leveraging multi-level contrastive learning loss, we fine-tune a language model (we select XLM-RoBERTa Vamvas and Sennrich (2023), see Appendix E) based on the human, human–LLM and LLM-generated texts, to force the model to reorganize the hidden space, pulling the embeddings within the same author families closer, and pushing the embeddings from different authors farther.
-We train an encoder that can represent text with distinguishable signals to discern authorship of text.
+We train an encoder that can represent text with distinguishable signals to discern the authorship of text.
 Here, we shift the focus to the detector, an encoder-based model that forms the core of FAID.
@@
   • If $x\in P_{2}$ ($x$ is generated by any LLM): Here, with $y_{\textit{LHS}}$ from $P_{3}$ and $y_{\textit{RHS}}$ from $P_{4}$ as defined above, the similarity $\mathrm{S}_{c}(x,y_{\textit{LHS}})$ is generally expected to be equal to $\mathrm{S}_{c}(x,y_{\textit{RHS}})$.
-Since $x$ can originate from any LLM, it doesn’t inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
+Since $x$ can originate from any LLM, it does not inherently possess a stronger connection to the specific LLM family in $P_{3}$ than to the broader human-LLM collaborations represented in $P_{4}$.
 This configuration aims to ensure that closeness in distribution corresponds to higher similarity after encoding, thus encouraging the model to discern fine-grained multi-level relations.
+Figure 2: Inference architecture: (a) embed the input text into embedding vector using the fine-tuned encoder, (b) use Fuzzy kNN to cluster, retrieving which cluster the input text belongs to (see more in Appendix F), (c) the stored vector database $\mathcal{VD}$ was created by saving all embeddings of texts in training and validation sets using the fine-tuned encoder.
+If the input text is unseen, we embed it and save it into a temporary vector database $\mathcal{VD}^{\prime}$, enhancing the generalization of the detector.
 4.2 Multi-level Contrastive Learning
-Given a dataset with $N$ samples, each sample is a text unit (paragraph/segment). The $i^{th}$ sample is denoted as $T_{i}$. Per sample, we assigned three-level labels indicating its source:
-  • $x_{i}\in\{0,1\}$: if $T_{i}$ is fully LLM-generated, $x_{i}=0$, otherwise $x_{i}=1$;
-  • $y_{i}\in\{0,1\}$: if $T_{i}$ is fully human-written, $y_{i}=0$, otherwise $y_{i}=1$;
+Given a dataset with $N$ examples, each example is a text unit (paragraph/segment). The $i^{th}$ record is denoted as $T_{i}$. Per record, we assigned three-level labels indicating its source:
+  • $x_{i}\in\{0,1\}$: if $T_{i}$ is a fully LLM-generated text, $x_{i}=0$, otherwise $x_{i}=1$;
+  • $y_{i}\in\{0,1\}$: if $T_{i}$ is a fully human-written text, $y_{i}=0$, otherwise $y_{i}=1$;
   • $z_{i}$: an indicator of a specific LLM family.
-The encoder $\phi(\cdot)$ represents the sample $T_{i}$ in a $d$-dimensional vector space $\mathbb{R}^{d}$.
-Then we calculate the cosine similarity between two samples $T_{i}$ and $T_{j}$, denoted by: $\sigma(i,j)=\mathrm{S}_{c}\left(\phi(T_{i}),\phi(T_{j})\right)$.
+The encoder $\phi(\cdot)$ represents the text $T_{i}$ in a $d$-dimensional vector space $\mathbb{R}^{d}$.
+Then we calculate the cosine similarity between two texts $T_{i}$ and $T_{j}$, denoted by: $\sigma(i,j)=\mathrm{S}_{c}\left(\phi(T_{i}),\phi(T_{j})\right)$.
 For LLM-generated text, the similarity between $T_{i}$ and another LLM-generated text $T_{j}$ is greater than that with a human-written or collaborative text $T_{k}$:
 $\sigma(i,j)>\sigma(i,k),\forall x_{i}=0,x_{i}=x_{j},x_{k}=1$ (2)
-If $x_{i}=0$, then the text sample is fully LLM-generated.
+If $x_{i}=0$, then the text is fully LLM-generated.
 For this case, we do not consider $y_{i}$ since LLM-generated text is considered as non-LLM-collaboration.
@@
 $\sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i}=y_{j},y_{i}\neq y_{k}$ (4)
-For cases where all samples are human–LLM collaborative, two texts created by the same LLM family tend to be more similar than the ones that involve contributions from different families. That is:
+For cases where all samples are human–LLM collaborative, two texts created by the same LLM family tend to be more similar than such that involve contributions from different LLM families. That is:
 $\sigma(i,j)>\sigma(i,k),\forall x_{i}=1,y_{i,j,k}=1,z_{i}=z_{j}\neq z_{k}$ (5)
@@
 $\mathcal{L}_{q}=-\log\frac{\exp\left(\sum_{k\in K^{+}}\frac{S(q,k)}{\tau}/N_{K^{+}}\right)}{\exp\left(\sum_{k\in K^{+}}\frac{S(q,k)}{\tau}/N_{K^{+}}\right)+\sum_{k\in K^{-}}\exp\left(\frac{S(q,k)}{\tau}\right)}$
-Different constraints result in varied sets of positive and negative samples.
+Different constraints yield different sets of positive and negative samples.
 Based on these sets, contrastive losses are computed at multiple levels.
@@
 $\begin{split}\mathcal{L}_{mcl}=&\sum\limits_{i=1}^{N}\left[x_{i}\left(\alpha\mathcal{L}_{q_{i},1}+\beta\mathcal{L}_{q_{i},2}\right)+\right.\\ &\left.\left(1-x_{i}\right)\left(\gamma\mathcal{L}_{q_{i},3}+\delta\mathcal{L}_{q_{i},4}+\zeta\mathcal{L}_{q_{i},5}\right)\right]\end{split}$ (8)
-Therefore, due to the last inequality in Eq (6) only specifying a case ($y_{i,j,k}=1$), and the other cases considering both values for $y$, so we have $\zeta=2\gamma=2\delta$.
+Due to the last inequality in Eq (6) only specifying a case ($y_{i,j,k}=1$), and the other cases considering both values for $y$, we have $\zeta=2\gamma=2\delta$.
 Also, to maintain the equilibrium, we need to keep $\alpha+\beta=\gamma+\delta+\zeta$.
@@
 As a result, it becomes more adept at recognizing variations in writing styles.
-This capability enhances its accuracy and strengthens its ability to generalize when detecting LLM-generated text.
+This capability enhances accuracy and strengthens generalizability when detecting LLM-generated text.
 4.3 Multi-Task Auxiliary Learning
@@
 Building on the previously described contrastive learning framework, we extend the encoder by attaching an MLP classifier at its output layer.
-This classifier is responsible for binary classification, determining whether a given text was written by a human or a LLM.
+This classifier performs binary classification, determining whether a given text was written by a human or an LLM.
 Let the probability of $i^{th}$ sample with label $x_{i}=0$ be $p_{i}$.
@@
 $\mathcal{L}=\mathcal{L}_{ce}+\mathcal{L}_{mcl}$ (10)
-Figure 2: Inference architecture: (a) embed the input text into embedding vector using the fine-tuned encoder, (b) use Fuzzy kNN to cluster, retrieving which cluster the input text belongs to (see more in Appendix F), (c) the stored vector database $\mathcal{VD}$ was created by saving all embeddings of texts in training and validation sets using the fine-tuned encoder.
-If the input text is unseen, we embed it and save it into a temporary vector database $\mathcal{VD}^{\prime}$, enhancing the generalization of the detector.
 4.4 Handling Unseen Data without Retraining
-Unseen data, whether from an unseen domain or an unfamiliar generator, remains a significant challenge even for state-of-the-art LLM-generated text detection methods, due to the increasing capabilities of LLMs.
+Unseen data, whether from an unseen domain or an unfamiliar generator, remains a significant challenge even for state-of-the-art LLM-generated text detection methods, as LLMs continue to improve.
 We tried to use the model classifier on its own, but we ended up using a vector database along with Fuzzy k-Nearest Neighbors, as illustrated in Figure 2.
-The results were written in Appendix F.
-Specifically, when dealing with unseen data, we use our model to embed these samples and add them to our existing vector database.
-Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without the need retraining.
+The results are given in Appendix F.
+Specifically, when dealing with unseen data, we use our model to embed these texts and add them to our existing vector database.
+Through careful parameter tuning, this approach enables our system to effectively handle newly encountered unseen data without retraining.
 5 Experiments
-In this section, we describe the datasets and baselines we used, followed by two experiments evaluating FAID: (i) classify a text as human, LLM, and human–LLM, and (ii) identify specific generators.
-Scenario Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
-In-domain,
-Known generator
+Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
 FAIDSet LLM-DetectAIve 94.34 94.45 93.79 94.10 0.1888 0.1107
@@
 FAID 96.73 97.61 98.05 97.80 0.4631 0.1806
-Unseen domain,
-Unseen generator
+Table 1: Performance with three labels. The best results are in bold and the second best are underlined.
+In this section, we describe the datasets and baselines we used, followed by two experiments evaluating FAID: (i) classify a text as human, LLM, and human–LLM, and (ii) identify specific generators.
+5.1 Datasets
+In addition to FAIDSet, for in-domain evaluation, we used two additional datasets:
+LLM-DetectAIve Abassy et al.
+(2024) encompasses various domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews.
+We augmented the original labels human-written and machine-generated using multiple LLMs to create a 485,405-example dataset with two new labels: (imachine-written then machine-humanized, and (iihuman-written then machine-polished.
+HART Bao et al.
+(2025) has 21,500 examples, including student essays, arXiv abstracts, story writing, and news articles.
+It covers four categories: human-written, AI-refined, AI-generated, and humanized AI-generated texts.
+The authors further expanded the dataset by creating additional instances with unbalanced label distributions.
+To evaluate the generalizability of FAID on unseen scenarios, we collected the following data:
+Unseen domain: We created a dataset consisting of 150 IELTS essays from Kaggle [^3] https://www.kaggle.com/datasets/mazlumi/ielts-writing-scored-essays-dataset, where all texts are human-written.
+We used these essays to generate human–LLM collaborative and LLM-generated texts with the same models with FAIDSet.
+Unseen generators: We selected 150 human-written abstracts from the FAIDSet test set and generated data for the remaining labels using three new LLM families: Qwen, Mistral, and Gemma.
+Unseen domain & generators: Based on the human-written IELTS essays above, we used the same LLM families as for the unseen generator test set to generate data for the LLM and the human–LLM labels.
+Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$ MSE $\downarrow$ MAE $\downarrow$
 Unseen domain LLM-DetectAIve 52.83 47.31 64.62 53.28 0.4733 0.4722
@@
 FAID 62.78 70.73 71.77 69.46 0.4514 0.4486
-Unseen
-generators
-LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
+Unseen generators LLM-DetectAIve 75.71 73.25 75.63 74.30 0.3714 0.2957
 T5-Sentinel 85.95 85.77 84.59 85.16 0.3648 0.2419
@@
 FAID 93.31 92.40 94.44 93.25 0.1691 0.1167
-Unseen domain
-and Unseen
-generators
-LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
+Unseen domains and Unseen generators LLM-DetectAIve 62.93 66.74 71.17 61.97 0.4479 0.3964
 T5-Sentinel 57.07 49.82 66.61 55.45 0.4314 0.4300
@@
 FAID 66.55 74.44 73.57 72.58 0.3939 0.3167
-Table 1: Performance with three labels.
-For in-domain and known generator datasets, we train the detectors on the training set and evaluate them on the test set.
-For unseen data evaluation, we use the detector trained on FAIDSet and evaluate on the unseen datasets.
-The best results are in bold and the second best are underlined.
-5.1 Datasets
-In addition to FAIDSet, for in-domain evaluation, we used two additional datasets:
-LLM-DetectAIve Abassy et al.
-(2024) encompasses various domains, including arXiv, Wikihow, Wikipedia, Reddit, student essays, and peer reviews.
-The original human-written and machine-generated texts were augmented by multiple LLMs to create a 236k-example dataset with two new labels: (imachine-written then machine-humanized, and (iihuman-written then machine-polished.
-HART Bao et al.
-(2025) with 21,500 examples including student essays, arXiv abstracts, story writing, and news articles.
-It contains four categories: human-written, AI-refined, AI-generated, and humanized AI-generated texts.
-The authors further expanded the dataset by creating additional instances with unbalanced label distributions.
-To evaluate the generalizability of FAID on unseen scenarios, we collected the following data:
-Unseen domain: We created a dataset consisting of 150 IELTS Writing essays from Kaggle ibrahimmazlum (2023), where all original texts are human-written.
-We used these essays to generate human–LLM collaborative and LLM-generated texts with the same models with FAIDSet.
-Unseen generators: We selected 150 human-written abstracts from the FAIDSet test set and generated data for the remaining labels using three new LLM families: Qwen, Mistral, and Gemma.
-Unseen domain & generators: Based on the human-written IELTS essays above, we used the same LLM families as for the unseen generator test set to generate data for the LLM and the human–LLM labels.
+Table 2: Performance with three labels with unseen data. We use the detector trained on FAIDSet and evaluate on the unseen datasets. The best results are in bold and the second best are underlined.
 Dataset Detector Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1-macro $\uparrow$
-FAIDSet LLM-DetectAIve 0.7596 0.7697 0.7690 0.7653
-T5-Sentinel 0.7568 0.7985 0.7840 0.7837
-SeqXGPT 0.6941 0.6802 0.6420 0.6603
-FAID 0.7964 0.8328 0.8352 0.8327
-LLM-DetectAIve LLM-DetectAIve 0.9049 0.9064 0.8352 0.8693
-T5-Sentinel 0.8154 0.8137 0.8009 0.8105
-SeqXGPT 0.8712 0.7890 0.7608 0.7441
-FAID 0.9089 0.8817 0.8672 0.8737
-HART LLM-DetectAIve 0.8900 0.8787 0.8674 0.8715
-T5-Sentinel 0.7852 0.7713 0.7834 0.7759
-SeqXGPT 0.6470 0.5582 0.4540 0.5075
-FAID 0.8996 0.9157 0.8648 0.8667
+FAIDSet LLM-DetectAIve 75.96 76.97 76.90 76.53
+T5-Sentinel 75.68 79.85 78.40 78.37
+SeqXGPT 69.41 68.02 64.20 66.03
+FAID 79.64 83.28 83.52 83.27
+LLM-DetectAIve LLM-DetectAIve 90.49 90.64 83.52 86.93
+T5-Sentinel 81.54 81.37 80.09 81.05
+SeqXGPT 87.12 78.90 76.08 74.41
+FAID 90.89 88.17 86.72 87.37
+HART LLM-DetectAIve 89.00 87.87 86.74 87.15
+T5-Sentinel 78.52 77.13 78.34 77.59
+SeqXGPT 64.70 55.82 45.40 50.75
+FAID 89.96 91.57 86.48 86.67
 Table 3: Accuracy of identifying generators: human, GPT, Gemini, Deepseek, and Llama. The best is in bold.
@@
 (2024) by fine-tuning a roberta-base sequence classification model.
-The input texts were tokenized using the RoBERTa tokenizer with a maximum sequence length of 256, and the model was trained for three-class detection.
+We tokenized the input texts using the RoBERTa tokenizer with a maximum sequence length of 256, and we trained the model for three-class detection.
 T5-Sentinel:
-We adapt the T5-Sentinel framework introduced by Chen et al.
+We adapted the T5-Sentinel framework introduced by Chen et al.
 (2023) for our three-class setup.
-Following the original configuration, we train using the AdamW optimizer (batch size 128, learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-5}$).
+Following the original configuration, we trained using the AdamW optimizer (batch size 128, learning rate $1\times 10^{-4}$, weight decay $5\times 10^{-5}$).
 This allows direct comparison with prior T5-based detectors under our experimental conditions.
 SeqXGPT:
-Following Wang et al.
-(2023) by modeling token-level likelihood patterns from LLM tokenizers for sentence-level detection, we adapted their method by updating the tokenizer models to align with our dataset’s label space.
-This adjustment ensures that the extracted token-level log-probability features correspond more accurately to the model types represented in our data.
-5.3 Human, LLM or Human-LLM?
-Table 1 shows the performance of FAID and three baselines in three evaluation settings.
-FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domains, (iii) unseen generators, and (iv) unseen domain & generators settings, followed by LLM-DetectAIve in (i) and (iv), and T5-Sentinel in (ii) and (iii).
-Although SeqXGPT aimed at extracting sequence-level features, it struggles with texts from advanced models, whose coherent and human-like writing styles reduce detectable distinctions.
-FAID improves generalization performance over unseen domains and generators compared to others.
+We adopted the method of Wang et al.
+(2023), who model token-level likelihood patterns from LLM tokenizers for sentence-level detection: we updated the tokenizer models to align with our dataset’s label space.
+This adjustment ensures that the extracted token-level log-probability features better reflect the model types in our data.
+5.3 Human-Only, LLM-Only, or Human-LLM?
+Tables 1 and 2 show the performance of FAID and three baselines in three evaluation settings.
+As shown in Table 1, FAID consistently achieves the best accuracy in (i) in-domain and known generators, (ii) unseen domains, (iii) unseen generators, and (iv) unseen domain & generators settings.
+It is followed by LLM-DetectAIve for (i) and (iv), and by T5-Sentinel for (ii) and (iii).
+Despite being designed to extract sequence-level features, SeqXGPT struggles with texts from advanced models, whose coherent, human-like writing styles reduce the detectable distinctions.
+FAID further improves generalization performance over unseen domains and generators compared to others as illustrated in Table 2.
 We can see that generalizing to (ii) unseen domains and (iv) unseen domain & generator remains challenging, with accuracy of 62.78% and 66.55%, respectively.
@@
 5.4 Identifying Different Generators
-We further examine the detector to identify specific AI generators seen during training.
-The goal of FAID is not only to detect AI involvement, but also to identify the specific LLM or family applied, treating them as distinct authors.
-As shown in Table 2, across the three datasets, FAID consistently achieves higher performance compared to other baselines in almost all metrics.
-LLM-DetectAIve achieves comparable results with our detector FAID on the test set of LLM-DetectAIve dataset, except that its precision is slightly lower.
-FAID’s high performance when dealing with text originating from diverse known generators within these datasets indicates that FAID learned unique writing patterns and features of different AI generators by leveraging multi-level contrastive learning.
+The goal of FAID is not only to detect whether AI was used to produce the target text, but also to identify the specific LLM family, treating the families as distinct authors.
+As shown in Table 3, FAID consistently achieves higher performance compared to other baselines in almost all evaluation measures.
+LLM-DetectAIve achieves results comparable to FAID on the test set of its dataset, except that its precision is slightly lower.
+FAID’s high performance when dealing with text from diverse known generators in these datasets indicates that it learned unique writing patterns and features of different generators by leveraging multi-level contrastive learning.
+Language Accuracy Precision Recall F1-macro
+English 96.41 96.02 95.59 95.77
+Vietnamese 94.42 95.60 94.22 94.76
+Table 4: Language-wise performance on FAIDSet.
+5.5 Language-wise Performance
+Table 4 presents the classification performance of FAID on each language of FAIDSet.
+The results show that FAID achieves consistently strong performance across both subsets.
+While performance on English is slightly higher, FAID maintains competitive accuracy and F1-score on Vietnamese, indicating that the model does not rely solely on high-resource language features.
+5.6 Generalizability to Unseen Human–LLM type of Collaboration
+To assess FAID’s robustness to unseen collaborative writing patterns between humans and LLMs, we conducted an additional experiment focusing on hybrid authorship styles that were not present in the original training dataset.
+We conducted manual quality control to enhance the reliability of our dataset and to ensure the robustness of the FAID model.
+A team of five annotators, all IT-majored, fluent in English, and aged 18–25, participated in the annotation process.
+Each annotator was assigned approximately 80 samples, covering all collaborative writing styles, for a total of 400 human-reviewed instances across various collaboration modes.
+During this manual revision stage, the annotators followed two key quality control principles:
+  1. 1.
+    Ensuring logical and informational consistency between the outputs and the original texts, where the output length was controlled to be 70–150% of the source length.
+  2. 2.
+    Improving quality through spelling correction, synonym replacement, and word refinement to ensure natural fluency and stylistic coherence.
+FAID was evaluated on this manually curated dataset without further fine-tuning.
+Despite these samples representing an unseen distribution, the model achieved a strong overall accuracy of 84.8%, precision of 82.8%, and recall of 85.0% across all collaboration categories.
+This suggests a strong generalization capability beyond the data distributions encountered during model development.
+It also highlights FAID’s sensitivity to fine-grained stylistic blending, where human revision only partially conceals the generative footprint of LLMs.
+5.7 Generalizability to Real-World Scenarios
+To further assess FAID’s generalizability beyond the controlled benchmark setting, we conducted an additional user study simulating realistic academic and professional writing scenarios.
+The goal was to evaluate whether FAID can maintain detection accuracy when applied to authentic, unconstrained text produced through real interactions with LLMs.
+A group of five volunteers with diverse academic backgrounds was instructed to engage with four popular AI systems: ChatGPT, Gemini, DeepSeek, and Llama 3.1 to generate text resembling authentic student or professional writing.
+The interactions were open-ended but were guided by three types of collaboration commonly encountered in two real-world use cases: writing a paper summary (as an abstract) and writing a passage for their own graduation thesis.Each participant was asked to generate five outputs per model, yielding a total of 200 real-world text samples.
+To ensure authenticity and diversity, the participants were encouraged to adjust the prompts iteratively, to include follow-up clarifications, and to perform light editing, mimicking realistic human–LLM co-writing behavior.
+FAID achieved strong performance, with overall accuracy of 88.5%, precision of 85.9%, and recall of 89.7%, indicating strong generalizability to unseen real-world scenarios despite being trained only in an in-domain setting.
 6 Conclusion and Future Work
-We presented FAIDSet, a multilingual fine-grained LLM-generated text detection dataset comprising 84k examples, and FAID, a framework designed to distinguish between human-written, LLM-generated, and human–LLM collaborative texts.
-FAID integrates multi-level contrastive learning and multi-task auxiliary objectives, treating LLM families as stylistic “authors,” which enables it to capture subtle linguistic and stylistic cues that generalize effectively across domains and evolving generative systems.
-Moreover, its Fuzzy KNN-based inference and training-free incremental adaptation contribute to strong robustness and adaptability to unseen data.
+We presented FAIDSet, a multi-domain, multilingual fine-grained LLM-generated text detection dataset comprising 83k examples, and FAID, a framework designed to distinguish between human-written, LLM-generated, and especially, human–LLM collaborative texts in practice.
+FAID integrates multi-level contrastive learning and multi-task auxiliary objectives, treating LLM families as stylistic “authors”, which enables it to capture subtle linguistic and stylistic cues that generalize effectively across domains and evolving generative systems.
+Moreover, its Fuzzy k-Nearest Neighbors-based inference and training-free incremental adaptation contribute to strong robustness and adaptability to unseen data.
 Our experiments demonstrated that FAID consistently outperforms competitive baselines across multiple datasets and settings.
 Its ability to detect nuanced collaborative writing and to adapt to emerging generative models highlights its potential for real-world deployment.
-In the future, we plan to extend FAIDSet to cover more languages, generators, and domains, particularly informal genres such as social media and student writing, to enhance cross-lingual and domain robustness further.
-We also aim to incorporate adversarial LLM-generated texts and more complex forms of human–LLM collaboration to capture the evolving dynamics of AI-assisted text creation.
-7 Limitations
+In future work, we plan to extend FAIDSet to cover more languages, generators, and domains, particularly low-resource languages and informal genres such as social media and student writing, to further enhance cross-lingual and domain robustness.
+We further plan to incorporate adversarially LLM-generated texts as well as more complex forms of human–LLM collaboration in order to better capture the evolving dynamics of AI-assisted text creation.
+Limitations
 While FAID demonstrates strong performance and generalization across various domains and LLMs, several limitations remain.
-First, although our dataset is multilingual and multi-domain, it is still limited in terms of low-resource languages and niche writing domains, which may affect performance in those contexts.
+First, although our dataset is multilingual and multi-domain, it remains limited in low-resource languages and niche writing domains, which may affect performance in those contexts.
 FAIDSet is synthetic by construction.
-The controlled generation enables causal-style analysis, but it under-represents the messy, tool-chain-specific edits, wich are seen in the wild.
+The controlled generation enables causal-style analysis, but it under-represents the messy, tool-chain-specific edits seen in the wild.
 We mitigate this with diverse prompt paraphrases, manual spot-checks, and generalization tests to held-out generators.
 Second, our framework is based on the observation that texts produced by LLMs from the same family share similar stylistic features.
-However, this may break down when a single text is influenced by multiple LLM systems – such as when a human uses different models for drafting, rewriting, and polishing a text.
+However, this may break down when a single text is influenced by multiple LLMs, e.g.,  when a human uses different models for drafting, rewriting, and polishing.
 In such cases, the resulting style may blend traits from multiple LLMs, making it more difficult to attribute authorship to a single LLM family or clearly distinguish collaboration boundaries.
-8 Ethics and Broader Impact
+Ethics and Broader Impact
 Data Collection and Licenses
 A primary ethical consideration is the data license.
-We reused existing datasets for our research: LLM-DetectAIve, HART, and IELTS Writing, which have been publicly released under clear licensing agreements.
-We adhere to the intended usage of these dataset licenses.
+We reused existing datasets for our research: LLM-DetectAIve, HART, and IELTS Writing, which have been publicly released with clear licenses and well-documented terms of use.
+We adhere to the intended usage of these datasets.
 Security Implications.
@@
 By spotting LLM-generated material, FAID helps preserve academic integrity, flag potential misconduct, and protect the genuine contributions of authors.
-More broadly, it supports efforts to prevent the misuse of generative technologies in areas such as credential falsification.
-Detecting LLM-generated content across different languages can be tricky, due to language’s grammar and style.
-By enabling robust, multilingual and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly.
-At the same time, it fosters critical digital literacy, giving everyone a clear understanding of both the strengths and limits of generative AI.
+More broadly, it supports efforts to prevent the misuse of generative technologies, such as credential falsification.
+Detecting LLM-generated content across different languages can be tricky, due to the language’s grammar and style.
+By enabling robust, multilingual, and multi-generator detection with accurate results, FAIDSet empowers people everywhere, especially in academic scenarios, to deploy AI responsibly.
+At the same time, it fosters critical digital literacy, giving everyone a clear understanding of both the strengths and the limits of generative AI.
 Responsible Use of AI-generated Text Detection.
 FAID is designed to enhance transparency in AI-assisted writing by enabling the fine-grained detection of AI involvement in text generation.
-While this has clear benefits for academic integrity, and content provenance, we acknowledge the potential for misuse.
+While this has clear benefits for academic integrity and content provenance, we acknowledge the potential for misuse.
 For instance, such tools could be used to unfairly penalize individuals in educational or professional settings based on incorrect or biased predictions.
@@
 Appendix A Some Current Datasets on AI-generated Texts
-We analyze many current datasets on AI-generated text and summarize them in Table 3.
-Many datasets contain fine-grained labels and multi-generators, but all of the analyzed datasets are monolingual – only in English.
-This problem prompts us to collect and formulate a new dataset for multilingual, multi-domain, and multilingual AI-generated text, enabling us to adapt the approach to real-world scenarios.
+We review existing datasets for AI-generated text detection and summarize their key characteristics in Table 5.
+While many of these datasets provide fine-grained labels and cover multiple text generators, all are monolingual and limited to English.
+This limitation highlights a critical gap in current resources and motivates the construction of a new multilingual, multi-domain, and multi-generator dataset for AI-generated text detection, which is essential for developing methods that generalize to real-world, cross-lingual scenarios.
 Dataset Languages Label Space Domains Generators Size
 MixSet
-Zhang et al. (2024c)
+Zhang et al. (2024)
 English
@@
 Table 5: English fine-grained AI-generated text detection datasets overview.
+Subset Human LLM Human–LLM
+collab.
+Train 14,176 12,076 32,091
+Validation 3,038 2,588 6,876
+Test 3,038 2,588 6,879
+Table 6: Number of examples per label in subsets in the FAIDSet dataset.
+Source Human Texts
+arXiv abstracts 2,000
+VJOL abstracts 2,195
+HUST theses (English) 4,898
+HUST theses (Vietnamese) 11,159
+Table 7: Statistics of human-written text’s origins in FAIDSet.
 Appendix B FAIDSet Statistics and Analysis
@@
 Our FAIDSet includes 83,350 examples, which are divided into three subsets: train, validation, and test, with the ratios shown in Table 6.
-The data set also comprises various sources of human-written texts, which are described in Table 5.
-Subset Human AI human–LLM collab.
-Train 14,176 12,076 32,091
-Validation 3,038 2,588 6,876
-Test 6,879 2,599 3,038
-Table 4: Number of examples per label in subsets in the FAIDSet dataset.
-Source No. Human texts
-arXiv abstracts 2,000
-VJOL abstracts 2,195
-HUST theses (English) 4,898
-HUST theses (Vietnamese) 11,164
-Table 5: Statistics of human-written text’s origins in FAIDSet.
+The dataset also comprises various sources of human-written text, as described in Table 7.
 B.2 Diverse Prompt Strategies
-To avoid biasing our generated corpora toward any single style or topic, we employ a broad set of prompt templates when synthesizing LLM-generated and human–LLM collaborative texts.
-By varying prompt structures, content domain, and complexity, we ensure that the resulting outputs cover a wide distribution of writing patterns, vocabulary, and rhetorical devices.
+In order to avoid biasing our generated corpora toward some style or topic, we use a broad set of prompt templates when synthesizing LLM-generated and human–LLM collaborative texts.
+By varying prompt structures, content domains, and complexity, we ensure that the resulting outputs cover a wide variety of writing patterns, vocabulary, and rhetorical devices.
 This diversity helps our detector generalize more effectively to real-world inputs.
-Concretely, we use five prompts for LLM-generated texts in Table 6 and several categories of human–LLM collaborative texts in Table 7, Table 8 and Table 9 consecutively:
+Concretely, we use five prompts for LLM-generated texts in Table 8 and several categories of human–LLM collaborative texts in Tables 9, 10, and 11, consecutively:
   • LLM-polished: Texts that a human initially wrote and then lightly refined by an LLM system to improve grammar, clarity, or fluency without altering the core content or intent.
@@
 Original abstract:   .
-Table 6: List of diverse prompt templates used to generate FAIDSet – Label: LLM-generated
+Table 8: List of diverse prompt templates used to generate FAIDSet – Label: LLM-generated.
 Student Thesis Paper Abstract
@@
 The original text:   .
-Table 7: List of diverse prompt templates used to generate FAIDSet – Label: LLM-polished
+Table 9: List of diverse prompt templates used to generate FAIDSet – Label: LLM-polished.
 Student Thesis Paper Abstract
@@
 Initial text:   .
-Table 8: List of diverse prompt templates used to generate FAIDSet – Label: LLM-continued
+Table 10: List of diverse prompt templates used to generate FAIDSet – Label: LLM-continued.
 Student Thesis Paper Abstract
@@
 The original text:
-Table 9: List of diverse prompt templates used to generate FAIDSet – Label: LLM-paraphrased
+Table 11: List of diverse prompt templates used to generate FAIDSet – Label: LLM-paraphrased.
 Appendix C Experimental Details and Real-World Use Cases
 C.1 Experimental Setup
-Unless noted otherwise, we use a batch size of 64, the AdamW optimizer with a learning rate of 2e52e-5 and weight decay of 1e41e-4, 50 epochs, and 2000 warm-up steps.
-For the fuzzy kNN component, we set top-K = 20.
-We use temperature = 0.7.
+Unless noted otherwise, we use a batch size of 64, the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and weight decay of $10^{-4}$, 50 epochs, and 2000 warm-up steps.
+For the fuzzy kNN component, we set top-K to 20, and we use a temperature of 0.7.
 C.2 Computational Cost
-All experiments run on a single NVIDIA A100 (40 GB); total wall-clock training time for FAID on FAIDSet is approximately 5 hours.
+We run all experiments on a single NVIDIA A100 (40 GB). The total wall-clock training time for FAID on FAIDSet is approximately 5 hours.
 In our setup using FAISS Douze et al.
-(2024) with CPU inference, the average query latency is approximately 5ms for FAIDSet with 84,000 embeddings on a standard server (Intel Xeon CPU, 64GB RAM).
-When performing with Fuzzy KNN, it takes some minor time to process, so the total average time for inference is 10ms.
+(2025) with CPU inference, the average query latency is approximately 5ms for FAIDSet with 83k embeddings on a standard server (Intel Xeon CPU, 64GB RAM).
+When performing with Fuzzy KNN, it takes some small period of time to process, and thus the total average time for inference is 10ms.
 C.3 Real-world Management of Vector Database
-In the case of FAIDSet, our entire training and validation embedding store requires approximately 200MB, which is easily handled by standard server disks.
+For FAIDSet, our entire training and validation embedding store requires approximately 200 MB, which is easily handled by standard server disks.
 For larger datasets, we believe the system remains scalable:
@@
   • Modern servers with hundreds of GB of disk are sufficient for storing large embedding banks, and the use of CPU-only inference (since Fuzzy KNN is GPU-free) makes the architecture cost-effective for deployment in resource-constrained environments.
-C.4 Generalizability of FAID on unseen human–LLM type of collaboration
-To further assess FAID’s robustness to unseen collaborative writing patterns between humans and LLM systems, we conducted an additional experiment focusing on hybrid authorship styles that were not present in the original training dataset.
-We performed a manual quality control process to enhance the reliability of our dataset and ensure the robustness of the FAID model.
-A team of five annotators, all IT-major university students fluent in English and aged 18-25, participated in the annotation process.
-Each annotator was assigned approximately 80 samples, covering all collaborative writing styles (LLM-generated, LLM-continued,LLM-paraphrased, and LLM-polished), resulting in a total of 400 human-reviewed instances representing various collaboration modes.
-During the manual revision stage, annotators followed two key quality control principles:
-  1. 1.
-    Ensuring logical and informational consistency between the output samples and original texts, where the output length was controlled to be between 70%–150% of the source length.
-  2. 2.
-    Improving linguistic quality through spelling correction, synonym replacement, and contextual word refinement to ensure natural fluency and stylistic coherence.
-FAID was evaluated on this manually curated dataset without any further fine-tuning.
-Despite these samples representing an unseen distribution, the model achieved a strong overall accuracy of 84.8%, precision of 82.8%, and recall of 85.0%, across all collaboration categories.
-This suggests a strong generalization capability beyond the data distributions encountered during model development.
-It also highlights FAID’s sensitivity to fine-grained stylistic blending, where human revision only partially conceals the generative footprint of LLMs.
-C.5 Generalizability of FAID on real-world scenarios.
-To further assess the generalizability of FAID beyond the controlled benchmark setting, we conducted an additional user study simulating realistic academic and professional writing scenarios.
-The goal was to evaluate whether FAID can maintain detection accuracy when applied to authentic, unconstrained text produced through real interactions with LLMs.
-A group of five volunteers with diverse academic backgrounds (same as in the Appendix C.4).
-Each participant was instructed to engage with four popular AI systems: ChatGPT, Gemini, DeepSeek, and Llama 3.1 on Together.AI to generate text resembling authentic student or professional writing tasks.
-The interactions were open-ended but guided by three types of prompts commonly encountered in real-world use cases: writing a paper summary (as an abstract) and writing a passage for their own graduation thesis.
-Each participant was asked to generate five outputs per model, yielding a total of 200 real-world text samples (5 participants $\times$ 4 models $\times$ 10 trials).
-To ensure authenticity and diversity, participants were encouraged to adjust prompts iteratively, include follow-up clarifications, and perform light editing, mimicking realistic human–LLM co-writing behavior.
-FAID achieved strong performance in this real-world setting, with an overall accuracy of 88.5%, precision of 85.9%, and recall of 89.7%, indicating that FAID still exhibits strong generalizability in uncontrolled real-world scenarios with in-domain settings.
 Appendix D Analysis on Similarity across Models and Model Families
-To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis from various perspectives, including N-gram distribution, text length patterns, and semantic embedding visualization.
-These analyses allow us to examine similarities both within the same model family and across different model families, leading to a robust understanding of AI "authorship" characteristics.
+To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis across multiple perspectives, including N-gram distributions, text length patterns, and semantic embedding visualizations.
+This allowed us to study the similarities within and across model families, leading to a robust understanding of AI “authorship” characteristics.
 D.1 Text Distribution between LLM Families
-We first examined the distribution of text length, measured in both word and character counts across outputs from five LLMs: Llama-3.3-70B-Instruct-Turbo Dubey et al.
-(2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Team et al.
-(2024).
-Using 2000 arXiv prompt seeds, each model generated a corresponding output, and their length distributions were plotted.
-From Figure 3, distinct patterns emerged between families.
-Gemini models consistently produced shorter and more compact outputs, while Llama and GPT models demonstrated more variance and a tendency toward longer completions.
-Despite using different versions (e.g., Gemini 2.0 vs.
-Gemini 1.5 Flash), the Gemini outputs remained closely grouped in terms of both word and character counts, suggesting a shared output generation strategy and stylistic consistency across the family.
-In contrast, Llama and GPT distributions showed greater separation and variability.
-This reinforces the hypothesis that text length behavior is not only model-dependent but also family-coherent, with Gemini models forming a distinct cluster.
+We analyzed the distribution of text length, measured by both word and character counts, across outputs from five LLMs: Llama-3.3-70B-Instruct-Turbo Dubey et al.
+(2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Gemini Team (2024).
+Using 2,000 arXiv prompt seeds, each model generated a single output, and we plotted the resulting length distributions.
+The results are shown in Figure 3, where we can observe that clear family-level patterns emerge.
+Gemini models consistently produce shorter, more compact outputs, whereas Llama and GPT models exhibit greater variance and a stronger tendency toward longer completions.
+Despite differences across versions (e.g., Gemini 2.0 vs.
+Gemini 1.5 Flash), Gemini outputs remain tightly clustered in both word and character counts, indicating a shared generation strategy and strong stylistic consistency within the family.
+In contrast, Llama and GPT distributions show greater dispersion and inter-model variability.
+This reinforces the hypothesis that text length is not only model-dependent but also family-coherent, with Gemini models forming a distinct cluster.
 D.2 Text Distribution between LLMs within the Same Family
@@
 D.3 Embedding Visualization and Semantic Cohesion
-To delve into deeper semantic alignment, we visualized embeddings generated by an unsupervised SimCSE XLM-RoBERTa-base model of generated texts from two model families: Gemini family (texts generated by Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash), and GPT-4o/4o-mini using PCA to reduce the high-dimensional embeddings into a space.
-As shown in Figure 4, Gemini model embeddings form tight, overlapping clusters, indicating a high degree of semantic cohesion among their outputs.
-This clustering remains consistent across both sample sizes of 2000 texts.
-In contrast, GPT-4o/4o-mini embeddings occupy a separate region of the space, with greater dispersion and less overlap with Gemini clusters.
-This visualization confirms that the Gemini family not only shares stylistic features but also maintains semantic coherence, distinguishing it from models of other families at a conceptual level.
+To explore semantic alignment in detail, we visualized the embeddings generated by an unsupervised SimCSE XLM-RoBERTa-base model on texts from two model families, Gemini and GPT, using PCA to project the high-dimensional embeddings into a lower-dimensional space for analysis.
+As shown in Figure 4, Gemini model embeddings form tight, overlapping clusters, indicating a high degree of semantic cohesion and internal consistency among their outputs.
+This clustering behavior remains stable across both sample sizes of 2,000 texts, suggesting that the observed patterns are not driven by sampling artifacts.
+In contrast, GPT-4o/4o-mini embeddings occupy a distinct region of the embedding space, exhibiting greater dispersion and noticeably less overlap with the Gemini clusters.
+Overall, this visualization confirms that the Gemini family not only shares stylistic features, but also demonstrates strong semantic coherence, effectively distinguishing it from models belonging to other families at a deeper conceptual level.
 D.4 Conclusion
@@
 Conversely, inter-family comparisons show clear separability, emphasizing the distinctiveness of each LLM family’s writing behavior.
-(a) Word length distribution across five LLMs.
-(b) Character length distribution across five LLMs.
+(a) word length distribution across five LLMs
+(b) character length distribution across five LLMs
 Figure 3: Text length distributions in words and characters across Llama-3.3, GPT-4o/4o-mini, Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash.
-(a) 2D embedding space.
-(b) 3D embedding space.
-Figure 4: Visualizations showing clustering behavior of Gemini model family (Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash) and GPT-4o/4o-mini using 2D and 3D embeddings with sample size of 2000 texts.
+(a) 2D embedding space
+(b) 3D embedding space
+Figure 4: Visualizations showing clustering behavior of Gemini model family (Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash) and GPT-4o/4o-mini using 2D and 3D embeddings with sample size of 2,000 texts.
 (a) Top 3-grams of Gemini 2.0 (500 samples).
-(b) Top 3-grams of Gemini 2.0 Flash (500 samples).
-(c) Top 3-grams of Gemini 1.5 Flash (500 samples).
+(b) top 3-grams of Gemini 2.0 Flash (500 samples)
+(c) top 3-grams of Gemini 1.5 Flash (500 samples)
 Figure 5: Top 20 most common trigrams from Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash using 500 sample prompts.
@@
 Table 12: Model selection on known vs. unseen generators. The best results in each column are in bold.
-To identify the best encoder for our classification task, we evaluated each candidate model on both known generators (FAIDSet testing data) and the new unseen-generator test set introduced in Section 3.
+In order to identify the best encoder for our classification task, we evaluated each candidate model on both known generators (in the FAIDSet test data) and the new unseen-generators test set, which is introduced in Section 3.
 Each transformer-based encoder was fine-tuned on FAIDSet training data and then used to predict labels on the two evaluation splits.
@@
 (2019) and e5‑base‑v2 Wang et al.
-(2022) achieved the most balanced trade‑off between classification accuracy and regression error (MSE, MAE), while Flan‑T5‑base Chung et al.
+(2024a) achieved the most balanced trade‑off between classification accuracy and regression error (MSE, MAE), while Flan‑T5‑base Chung et al.
 (2022) lagged slightly in F1‑macro.
@@
 Multilingual variants.
-Next, we experimented with XLM‑RoBERTa‑base Conneau et al.
-(2019) and Multilingual‑e5‑base Wang et al.
-(2024a).
-Both models leverage cross-lingual pretraining, which, in our scenario, appears to improve the representation of diverse linguistic patterns present in FAIDSet.
-In particular, XLM‑RoBERTa‑base delivered a substantial boost in all metrics, suggesting that its multilingual training enhances generalization even on primarily monolingual inputs.
+We next evaluated XLM-RoBERTa-base Conneau et al.
+(2020) and Multilingual-e5-base Wang et al.
+(2024b).
+Both models benefit from cross-lingual pretraining, which in our setting improves the representation of the diverse linguistic patterns present in FAIDSet.
+Notably, XLM-RoBERTa-base yields a substantial improvement across all evaluation metrics, indicating that its multilingual training enhances generalization even when applied to predominantly monolingual inputs.
 Contrastive learning with SimCSE.
-Finally, we integrated contrastive learning via SimCSE Gao et al.
-(2021) to further refine the sentence embeddings.
-We tested both supervised (trained on NLI data) and unsupervised (trained on Wikipedia corpus) SimCSE variants applied to RoBERTa‑base.
-The unsupervised version outperformed its supervised counterpart, confirming prior findings that unsupervised SimCSE often yields stronger semantic encoders for downstream tasks.
-Encouraged by this, we adopted the unsupervised SimCSE approach on XLM‑RoBERTa‑base Vamvas and Sennrich (2023), which achieved the highest accuracy and lowest error rates across all configurations.
+Finally, we incorporated contrastive learning via SimCSE Gao et al.
+(2021) to refine sentence embeddings.
+We evaluated supervised (trained on NLI data) and unsupervised (trained on the Wikipedia corpus) SimCSE variants applied to RoBERTa-base.
+The unsupervised variant outperformed its supervised counterpart, aligning with prior findings that unsupervised SimCSE produces stronger semantic encoders.
+Based on these results, we applied unsupervised SimCSE to XLM-RoBERTa-base Vamvas and Sennrich (2023), achieving the highest accuracy and lowest error rates.
 Based on these experiments, we selected the unsupervised SimCSE XLM‑RoBERTa‑base model for our final system.
@@
 F.1 The Need to Use a Vector Database
-We first used the trained model to classify the text and observed a significant performance decline from 90% when evaluating on unseen domains or text generated by novel models.
-This drop reflects the model’s limited ability to generalize beyond the distribution of its original training data.
-To address this, we propose integrating a vector database that stores dense embeddings of all collected examples, both training and incoming unlabeled data.
-By indexing and retrieving semantically similar examples at inference time, the vector database serves as a flexible, scalable memory that bridges gaps between training and test distributions, enhancing the classifier’s resilience to domain shifts and unseen generators.
+We first applied the trained model to classify the input text and observed a substantial performance drop of 15–30% when evaluated on unseen data.
+This degradation underscores the model’s limited ability to generalize beyond its training distribution and its sensitivity to distributional shifts.
+To address this issue, we decided to integrate a vector database that stores dense embeddings of all examples, including both labeled instances and unlabeled data encountered during inference.
+By indexing and retrieving semantically similar examples during inference, the vector database serves as a flexible, scalable memory module that helps bridge the gap between the training and test distributions.
+This retrieval-based mechanism enhances the classifier’s robustness to domain shifts and unseen generators by grounding predictions in stylistically and semantically related examples.
 Specifically:
@@
 F.2 Clustering Algorithm Selection
-To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies on the encoded vectors stored in our vector database.
+To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies in our vector database.
 Each algorithm was tasked with grouping text samples into human‐written, AI‐generated, and human–LLM collaborative categories, using both known‐generator data (held out from training) and entirely unseen generator data.
-Performance was measured in terms of accuracy and F1-macro score.
-We encoded each text sample using the penultimate layer of our pre-trained sequence-classification model, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three text-origin classes.
-From the results in Table 11, we draw some key findings:
-  • Traditional algorithms show reasonable performance on held‐out known generators but degrade notably on unseen generators.
-  • Fuzzy C‐Means leverages membership degrees to handle overlapping distributions, improving both accuracy and F1-macro by about 4% over k‐Means, with smaller degradation on unseen data.
-  • Fuzzy KNN combines local neighbor information with fuzzy membership, achieving the best overall performance: 95.2% accuracy and 95.1% F1-macro on known generators, and 93.3% accuracy and 93.3% F1-macro on unseen generators.
-Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our whole architecture.
+For evaluation, we used accuracy and F1-macro score.
+We encoded each example using the penultimate layer of our classification model, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three classes.
+The results are shown in Table 13, where we can see:
+  • Traditional algorithms show reasonable performance on held‐out known generators, but degrade notably on unseen generators.
+  • Fuzzy C‐Means leverages membership degrees to handle overlapping distributions, improving both measures by 4% over k‐Means, with smaller degradation on unseen data.
+  • Fuzzy KNN combines local neighbor information with fuzzy membership, achieving the best overall performance.
+Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our overall architecture.